### v3

# EvalNet: A Practical Toolchain for Generation and Analysis of Extreme-Scale Interconnects

Abstract.
The diversity of communication paths in a network—especially non-minimal paths—is a key enabler of performance at extreme scales. We present EvalNet, a toolchain for scalable generation and analysis over 25 important network topologies, such as Slim Fly, PolarFly, and Orthogonal Fat Trees, with a strong focus on path diversity metrics. EvalNet provides an extensive and fine-grained analysis of shortest and non-shortest paths, including their multiplicities, lengths, and interference. It supports exact measurement and visualization of bandwidth and throughput between every router pair, enabling unprecedented insight into routing potential. EvalNet also includes detailed models for construction cost and power consumption, and interfaces seamlessly with established simulators, which we tune to support large-scale evaluations on low-cost hardware. Using EvalNet, we deliver the widest and most comprehensive path diversity study to date, demonstrating how path diversity underpins throughput and scalability, and facilitating progress towards new frontiers in extreme-scale network design.
copyright: noneccs: Networks Network performance evaluationccs: Networks Network performance analysisccs: Networks Network experimentationccs: Networks Data center networksccs: Networks Packet-switching networks
1. INTRODUCTION
The interconnection network plays an important role in today’s large-scale computing systems (Besta and Hoefler, 2014; Niranjan Mysore et al., 2009; Handley et al., 2017; Valadarsky et al., 2015; Besta et al., 2021; Di Girolamo et al., 2019; Copik et al., 2021; Hoefler et al., 2020; Di Girolamo et al., 2022). Large networks with tens of thousands of nodes are deployed in warehouse-sized HPC and data centers (Fu et al., 2016; Liao et al., 2014). Future supercomputers as well as mega data centers will require even larger scales with hundreds of thousands of servers.
A modern network topology must satisfy many stringent requirements related to different aspects of the interconnect design, such as high performance or low construction cost. However, the evaluation of all these aspects is challenging. First, there are many networks that must be considered for comparison when designing a new topology, for example Fat Tree (Leiserson et al., 1996), Hypercube (Dally and Towles, 2003), Torus (Dally and Towles, 2003), Dragonfly (Kim et al., 2008), Slim Fly (Besta and Hoefler, 2014), Xpander (Ahn et al., 2009), or Jellyfish (Singla et al., 2012) – to name only a few. On top of that, many of these baseline networks, for example Xpander or Orthogonal Fat Trees, do not have straightforward construction algorithms, making it challenging and time-consuming to compare to these baselines.
Similarly, there are many performance measures that should be considered for overall evaluation, and these measures often require different tools. A core challenge in evaluating interconnects is the analysis of path diversity, which has emerged as a critical enabler of performance and scalability in modern networks (Besta et al., 2020). As networks grow to extreme sizes, there is a growing need for fine-grained and scalable tools that can assess not only the existence of diverse paths but also their structure, interference, and utility for multipath routing. In parallel, several complementary performance metrics must also be considered, such as throughput and latency. Throughput is ideally evaluated using packet-level simulators, which offer high accuracy but often do not scale to large topologies. Flow-based simulations scale more easily, but they tend to produce less accurate predictions (Casanova et al., 2008). Furthermore, additional aspects like construction cost and power consumption must be analyzed. Unfortunately, the tools supporting these various metrics often rely on different formats for specifying networks, which significantly complicates integrated and efficient evaluation.
To address all these issues, we propose EvalNet: a publicly available toolchain that facilitates the integrated and comprehensive evaluation of networks. EvalNet consists of several modules, some of which were developed from scratch, while others are interfaces to existing tools which facilitate seamless operation. First, EvalNet enables generation of a very large number of network topologies. We cover the state-of-the-art class of low-diameter topologies developed by optimizing the network structure towards the Moore Bound (Slim Fly (Besta and Hoefler, 2014), PolarFly (Lakhotia et al., 2022), BundleFly (Lei et al., 2020), Xpander (Valadarsky et al., 2015)), Dragonfly topologies (fully-connected Dragonfly (Kim et al., 2008), Cascade Dragonfly (Faanes et al., 2012)), random networks (Jellyfish (Singla et al., 2012)), torus related topologies (Mesh, Torus, Hypercube (Dally and Towles, 2003), Express Mesh (Jain et al., 2017), HyperX (Ahn et al., 2009), Flattened Butterfly (Kim et al., 2007), Tofu (Ajima et al., 2018, 2012)), indirect topologies (Fat Tree (Leiserson et al., 1996), Multi-Layer Full-Mesh (Kathareios et al., 2015), Orthogonal Fat Tree (Kathareios et al., 2015), eXtended Generalized Fat Tree (Ohring et al., 1995), k-ary n-tree (Petrini and Vanneschi, 1997)), and Kautz networks (Undirected Kautz (Li et al., 2004), Arrangement Network (Day and Tripathi, 1992)).
Figure 1. An overview of EvalNet.
[…]
[Figure is fully described in the text.]Figure is fully described in the text.
Second, we observe that the high performance of network workloads is becoming increasingly dependent on effectively harnessing various forms of path diversity, particularly through non-minimal multipathing (Besta et al., 2020). This is especially critical in state-of-the-art low-diameter topologies, such as Slim Fly and PolarFly, which typically offer limited shortest-path diversity but compensate with a rich structure of non-minimal paths. In such networks, the ability to balance load and avoid congestion hinges on the capacity to exploit this non-shortest-path structure. To address this need, EvalNet provides a comprehensive suite of fine-grained tools for analyzing path diversity. It supports the computation of shortest and non-shortest path multiplicities, detailed distributions of path lengths, and interference-aware metrics that capture how routes interact with each other. EvalNet also includes visualization and measurement tools that expose the exact bandwidth and multipathing potential between every pair of routers, enabling both coarse-grained insights and fine-grained optimization of routing and scheduling schemes (Lakhotia et al., 2023). All path diversity analyses are designed to scale to very large topologies, supporting networks with more than 100,000 servers, thus making them practical for evaluating future-generation systems.
Moreover, EvalNet seamlessly interfaces with different tools for network simulation. This includes the flow-level evaluation tool TopoBench (Jyothi et al., 2016) that enables approximating throughput for arbitrary traffic patterns and comes with a specially designed worst-case scenario that puts particularly large stress on the interconnect. EvalNet also harnesses two complementary packet-level simulators (htsim (Raiciu et al., 2010) and Booksim2 (Jiang et al., 2013)), which target, respectively, lossy and lossless architectures [^1] EvalNet does not compete with existing simulators as it does not introduce a new simulator. Instead, it tunes two of these simulators and harnesses them within its full comprehensive pipeline.. We tune these simulators to support large-scale evaluations on low-cost hardware, achieving simulations of more than 1 million endpoint networks on a commodity PC. Finally, EvalNet also comes with a construction cost and power consumption assessment module. The user can parametrize this module with arbitrary cost and power models for both routers and cables, considering arbitrary vendors and architectures such as Ethernet or InfiniBand.
Finally, we use EvalNet to conduct the widest-ever comparative analysis of interconnect topologies, focusing on the path diversity measures, but also combining them with classical metrics that include latency, global bandwidth, construction cost, and power consumption. We investigate path diversity in-depth because other measures (latency, bandwidth under flow/packet simulations, fault-tolerance) have already been extensively researched. However, path diversity has only been analyzed in a limited scope. We provide a total of 55 rich plots for path diversity metrics, combined with observations, analyses, and insights that could be used to develop better routing strategies.
2. EVALNET TOOLCHAIN
We illustrate the overall structure of EvalNet in Figure 1. First, the user selects a topology to be generated; it can be any of the networks described in Section 3; the user can also provide a script generating their own interconnect. The user can then use a module for analyzing the path diversity of a network, which can subsequently be used to enhance routing schemes in simulations (an overview is in Section 4.1, details and evaluation are in Section 5). The generated networks can also be evaluated using either flow or packet-based simulations (overview is in Section 4.2, details and evaluation are in Section 6 and 7, respectively). These simulations are conducted by external tools (TopoBench (Jyothi et al., 2016) for flow-based and htsim (Raiciu et al., 2010) as well as Booksim2 (Jiang et al., 2013) for packet-based simulations of lossy and lossless networks, respectively); we provide input/output conversion routines to integrate these tools seamlessly. Finally, the user can use our cost and power models to assess the construction cost or static power consumption (Section 4.3).
We now proceed to outline these factors (Sections 34) and then detail each of them and provide extensive evaluations (Sections 57). We will finalize with a comprehensive summary comparing all considered interconnects and measures (Section 8).
3. RELEVANT NETWORK TOPOLOGIES
We first describe network topologies included in EvalNet – see also Figure 2. The most important symbols are listed in Table 1.
$V,E$ Sets of vertices/edges (routers/links, $V=\{0,\dots,N_{r}-1\}$).
$N,N_{r}$endpoints and #routers in the network ($N_{r}=|V|$).
$p,k^{\prime}$endpoints attached to a router, #channels to other routers.
$k$ Router radix ($k=k^{\prime}+p$).
$D,d$ Network diameter and the average path length.
$c_{l}(A,B)$ Count of (at most $l$-hop) disjoint paths between router sets $A$, $B$.
$c_{\text{min}}(s,t),l_{\text{min}}(s,t)$ Diversity and lengths of minimal paths between routers $s$ and $t$.
$I_{ac,bd}$ Path interference between pairs of routers $a,b$ and $c,d$.
$\lambda,v$ Flow arrival rate [flows/s] and the flow volume [bytes].
Table 1. The most important symbols used in this work.
Figure 2. An overview of network topologies offered by default in EvalNet. New topologies can be straightforwardly added to the pipeline.
[…]
[Figure is fully described in the text.]Figure is fully described in the text.
3.1. Network Model
We model an interconnection network as an undirected graph $G=(V,E)$; $V$ and $E$ are sets of vertices and edges. A vertex models a router [^2] We abstract away HW details and use “router” for both L2 switches and L3 routers. ($|V|=N_{r}$). An edge models a full-duplex inter-router physical link. Servers (endpoints) are modeled implicitly. There are $N$ servers in total, and $p$ endpoints are attached to each router (concentration). There are $k^{\prime}$ channels between any two routers (network radix). $k=p+k^{\prime}$ is the total router radix. The diameter is $D$ while the average path length is $d$. $\lambda$ and $v$ are the flow arrival rate [flows/s] and the flow volume [bytes] of a used workload.
While EvalNet supports arbitrary configurations of $p$, $k$, and $k^{\prime}$, in the experiments, we usually set $p=\frac{k^{\prime}}{d}$, as it maximizes throughput while minimizing congestion and network cost (assuming a random uniform traffic pattern) (Besta et al., 2020). We also usually select network radix $k^{\prime}$ and router count $N_{r}$ so that, for a fixed $N$, the compared topologies use similar amounts of networking hardware and thus have similar construction costs. We assume router-based topologies hereafter, i.e., no direct channels between any two endpoints.
3.2. Moore Bound Topologies
A recent class of low-diameter topologies explicitly optimizes the network structure towards the Moore Bound (Besta and Hoefler, 2014). This maximizes the number of routers and endpoints for a fixed diameter and radix, minimizing cost and power consumption. These topologies are important as they have achieved new frontiers in performance, cost, and power consumption.
Slim Fly (SF) (Besta and Hoefler, 2014), PolarFly (PF) (Lakhotia et al., 2022), and PolarStar (Lakhotia et al., 2024) belong to this class, they are based on the underlying McKay-Miller-Širáň and Erdős-Rényi polar graphs, respectively. Spectralfly (SpF) (Young et al., 2022) is based on the Ramanujan graph construction of Lubotzky, Phillips, and Sarnak (LPS). Moreover, a $\tau$-regular Xpander (XP) (Valadarsky et al., 2015) constructions starts with a complete graph of $\tau+1$ vertices, and is then repeatedly “lifted” ($\frac{N_{r}}{\tau+1}$ times) to obtain the final graph with good expansion properties. Finally, BundleFly (BF) (Lei et al., 2020) is using SF together with a Paley graph. The routers in both graphs are linked via a multi-star product based on a bijection.
3.3. Dragonfly Topologies
The Fully-Connected Dragonfly (DF) (Kim et al., 2008) is comprised of $g$ groups of nodes internally interconnected in an all-to-all fashion. The topology among groups is an all-to-all, too. A Cascade Dragonfly (C-DF) (Faanes et al., 2012) alters the original Dragonfly by assuming a 16x6 2D HyperX for intra-group and an all-to-all graph for inter-group connectivity. Megafly (MF) (Flajslik et al., 2018) is another Dragonfly variant. These networks are relevant as they come with good performance and cost, albeit achieving worse values for these metrics than low-diameter networks, but have been found to be easier to cable. They also offer relatively easy scheduling: their modular structure facilitates achieving data locality for workloads (Tate et al., 2014).
3.4. Random Topologies
Fully random networks are infeasible due to the unbounded radix. The Jellyfish (Singla et al., 2012) is a random graph with a pre-selected router radix $k$. We support these networks as they come with advantageous performance, cost, and power consumption (Singla et al., 2012; Besta et al., 2020).
Figure 3. An overview of evaluation measures of network topologies offered by default in EvalNet. New measures can be straightforwardly added to the pipeline.
[…]
[Overview of evaluation measures for network topologies.]The figure provides an overview of the structure of the Evaluating Network Topologies chapter as well as the considered metrics. The chapter consists of path diversity measures (4.1), cost and power (4.3) as well as throughput and latency (4.2) subsections. The last subsection considers throughput (saturation for a given workload), maximum achievable throughput (MAT), network latency and average path length. The second subsection looks at the construction cost and power consumption. The path diversity subsection investigates counts for minimal and non-minimal paths as well as path interference and total network load (TNL). For the minimal and non-minial paths specifically lengths and multiplicity are considered.
3.5. Torus Related Topologies
We consider traditional Mesh (MH), Torus (TR), and Hypercube (HC) topologies. We also incorporate their more recent enhancements: HyperX (HX) (Ahn et al., 2009) (each router in $n$-D grid connects to all other routers within the same dimension), Express Meshes (EM) (Jain et al., 2017) (cut-down versions of the HyperX), Flattened Butterfly (FBF) (Kim et al., 2007) (a 2D variation of HyperX), and Tofu (TF) (Ajima et al., 2018, 2012) (a 6D torus variant with inner 3D mesh/tori networks embedded as virtual routers). These networks have been widely deployed across different settings and thus are considered in EvalNet.
3.6. Indirect Topologies
In indirect networks, some routers attach endpoints while other routers only forward packets. A Fat Tree (FT) (Leiserson et al., 1996) is a binary tree, with either higher bandwidth links towards the root or multiple parallel links towards the root to avoid communication bottlenecks. It is one of the most widely researched and used network topologies (Katta et al., 2016b; Vanini et al., 2017; Alizadeh et al., 2014; Katta et al., 2016a; He et al., 2015; Zhang et al., 2017; Huang et al., 2020; Alasmar et al., 2018; Montazeri et al., 2018; Hwang et al., 2014; Lu, 2016; Vamanan et al., 2012; Alizadeh et al., 2013; Bai et al., 2014; Raiciu et al., 2011; Handley et al., 2017; Zhuo et al., 2016; He et al., 2016; Alizadeh et al., 2010; Cardwell et al., 2017; Mittal et al., 2015; Li and Pan, 2013; Perry et al., 2014; Olteanu and Raiciu, 2016; Park et al., 2019; Benet et al., 2018; Lu et al., 2018; Widmer et al., 2001; Callegari et al., 2014). The k-ary n-tree (AT) (Petrini and Vanneschi, 1997) avoids infeasible high-radix roots of FT by using the same number of routers per tree-level and connecting them with increasing strides towards the top. An eXtended Generalized Fat Tree (xGFT) (Ohring et al., 1995) is a k-ary n-tree variant which allows reducing deployment cost. Orthogonal Fat-Trees (OFT) (Kathareios et al., 2015) are based on k-ary 2-trees; the wiring pattern between the two levels is modified to increase scalability. Finally, a Multi-Layer Full-Mesh (MLFM) (Kathareios et al., 2015) consists of multiple 2D meshes (incl. diagonal connections) stacked together, but instead of direct links it uses global routers to link meshes. EvalNet also supports indirect networks as they offer very high bandwidth; FTs are also widely deployed.
3.7. Kautz Topologies
Both Undirected Kautz (KA) (Li et al., 2004) and Arrangement Network (Day and Tripathi, 1992) have strings assigned to each router. Then, links between routers are determined by similarities of the corresponding router strings. Kautz networks offer high resilience properties and offer an interesting alternative to other classes of interconnects.
3.8. Other Topologies
EvalNet can be easily extended to model and evaluate interconnects from other domains, such as on-chip networks between different cores (Ahn et al., 2015; Besta et al., 2018), chiplets (Iff et al., 2023a, 2025, b), or FPGA interconnects (Cai et al., 2015), due to its modular and flexible architecture. The toolchain’s seamless integration with external simulators for packet-level and flow-level analysis ensures that new topologies, such as those designed for heterogeneous (Chang et al., 2012) or spatial computing environments (Gianinazzi et al., 2023; Baumann et al., 2024), can be evaluated with minimal effort, making EvalNet a versatile tool for broader interconnect research.
4. EVALUATING NETWORK TOPOLOGIES
We next describe the considered measures used to evaluate a network topology. We analyzed a large body of related work to provide a comprehensive overview (Besta et al., 2021; Kim et al., 2008, 2007; Besta and Hoefler, 2014; Prisacari et al., 2013b; Yuan et al., 2014, 2013; Prisacari et al., 2014a, b; Kathareios et al., 2015; Prisacari et al., 2015; Chen et al., 2016; Prisacari et al., 2013a; Karacali et al., 2018; Sehery and Clancy, 2017; Kassing et al., 2017; Dally and Towles, 2003). We distinguish five classes of measures, used to analyze five aspects of a network: path diversity, bandwidth, latency, construction cost, and power consumption. All these metrics are also listed in Figure 3.
4.1. Path Diversity Measures
Recent studies in high-performance interconnect design indicate that harnessing different forms of path diversity, especially non-minimal multipathing, is a key performance enabler, especially for state-of-the-art low-diameter networks such as SF or PF. This is because these networks – while ensuring low latency, high throughput, as well as low construction cost and power consumption – have only one shortest path between almost all router pairs (Besta et al., 2020, 2021). Hence, one cannot use traditional load-balancing mechanisms such as equal-cost multi-path routing (ECMP) (Hopps, 2000) for more performance, and must resort to multipath routing over non-minimal paths (Besta et al., 2020).
To address this, EvalNet supports analyzing all recently proposed measures for path diversity. First, it enables computing counts of both minimal and non-minimal paths between any two routers. These two measures illustrate the potential for multipathing over both traditional and recent networks. Moreover, we use the path interference (PI) measure, which illustrates how multipathing changes when considering interactions between different pairs of routers (i.e., some parts of a given network may be more or less advantageous for multipathing than other parts of the same network). This enables deriving detailed insights into path diversity in the local network structure, facilitating the fine-grained design of routing protocols. Finally, EvalNet enables computing the total network load (TNL), which is an upper bound on the number of flows that a network can maintain without congestion; TNL constitutes the maximum supply of path diversity offered by a given topology. We define all these measures in more detail in Section 5 while analyzing their values for different topologies.
Many of the above measures can be analyzed at different levels of detail. They can be summarized with histograms, or they can be computed and plotted separately for each individual router pair. EvalNet supports both options.
Figure 4. (left) Histograms of shortest path lengths $l_{\text{min}}$. (right) Histograms of shortest path multiplicities $n_{\text{min}}$. Considered networks: DF, SF, and two HX variants with $D\in\{2,3\}$.
[…]
[Histograms of shortest path lengths and multiplicities]The figure shows in two subfigures the histograms of the shortest path lengths and multiplicities for four topologies: Dragonfly, Slim Fly and HyperX (diameter of 2 and 3). Four different server sizes are illustrated: 1.000, 10.000, 100.000 and 1.000.000. The shortest path lengths facilitate reasoning about the expected latency of a topology with shortest-path minimum routing, for different router pairs. More router pairs in Slim Fly are connected with a 2-hop shortest path, than in HyperX (diameter of 2), for all topology sizes. This motivates more focus on making best-possible routing decisions over 2-hop paths in Slim Fly; 1-hop paths are negligible (especially for large networks). The shortest path multiplicities enable taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing. Nearly all router pairs in Slim Fly have only 1 shortest path between them, making non-minimal multipathing crucial. Most router pairs in HyperX have many shortest paths between them, and this ratio is growing with the number of servers, making ECMP a good option.
4.2. Throughput & Latency Measures
In EvalNet, we offer seamless interfaces to established tools, facilitating assessing throughput using packet-based simulations. We define throughput to be the normalized rate of injecting input packets for which the network saturates, i.e., the average latency of packets becomes higher than a pre-specified threshold. Moreover, we also consider Maximum Achievable Throughput (MAT), i.e., an approximation of throughput obtained using flow-based simulations conducted with TopoBench. MAT is defined as the ratio of the flow that can be forwarded between each pair of endpoints, to the demanded amount of flow for this pair.
Network latency is also obtained using packet-based simulations. The network latency is defined as the latency achieved for a very low injection rate that is close to zero. Thus, latency measures the end-to-end communication delay without congestion effects in the network. Moreover, the interfaced simulators enable delivering the average path length, i.e., the average number of router-router links traversed between any two communicated endpoints. This measure is useful because it is related to the network topological structure, i.e., one can see more clearly (than with latency) the impact of a routing design decision on the number of links that must be traversed.
4.3. Construction Cost & Power Consumption
EvalNet enables assessing the network construction cost. The user can plug in different cable and router cost models, coming from arbitrary vendors such as Ethernet or InfiniBand, and incorporating different rack placement. EvalNet also offers a model assessing the base power consumption required to power the network. This enables rudimentary comparisons between different baselines and preliminary insights into power consumption; detailed power models are notoriously hard to devise and vendors resort to their in-house infrastructures.
5. PATH DIVERSITY ANALYSIS
We present path diversity measures and conduct analyses in EvalNet, obtaining interesting insights about various interconnects. These measures are defined on pairs of routers (Besta et al., 2020). Extended results are provided in the Appendix. We start with basic definitions shared by all measures.
Definition 5.1.
The neighborhood $h(A)$ of a set of vertices $A\subseteq V$ contains all the vertices adjacent to vertices in $A$, i.e., $h(A)=\left\{t\in V:\exists_{s\in A}(s,t)\in E\right\}.$
Intuitively, $h(A)$ are all the routers connected to routers in set $A$.
Definition 5.2.
The $l$-step neighborhood $h^{l}(A)$ (or a transitive hull) of $A\subseteq V$ contains all the vertices reachable by a path of length $l$ from some vertex in $A$. Hence, $h^{l}(A)=\underbrace{h(\cdots h(}_{l\text{\ times}}A)\cdots).$
Intuitively, $h^{l}(A)$ are all the routers reachable with $l$ hops from routers in set $A$.
Definition 5.3.
For $s,t\in V$, the number of length-$l$ paths $n_{l}(s,t)$ from $s$ to $t$ is recursively defined as $n_{l}(s,t)=\sum_{u\in V}n_{l-1}(s,u)\cdot n_{1}(u,t),$ $n_{1}(s,t)=1\text{ iff }(s,t)\in E,\text{ else }0.$
Figure 5. The analysis of non-shortest path diversity and path interference.
[…]
[Histograms of multiplicities of non-shortest paths and path interferences for shortest & non-shortest paths as well as the path diversity between all router pairs and 2D histograms of path interferences for shortest & non-shortest paths]The figure analyses non-shortest path diversity and path interference in four subfigures. The compared topologies are Dragonfly, SlimFly and HyperX (diameter of 2 and 3). The multiplicities of non-shortest paths facilitate reasoning about the expected latency of a topology with non-shortest path routing, for different router pairs, and offers hints on routing design. HyperX with a diameter of 2 has very high path diversity that is also ’stable’ (20 for all router pairs) for all higher path lengths (l=3,4,5). This facilitates the design of complex adaptive routing that routes around hotspots, uses complex paths, et cetera. The analysis of path interferences for shortest & non-shortest paths facilitates reasoning about the performance of both minimal and non-minimal routing, and of scheduling of potential processes to different servers. Dragonfly, HyperX with a diameter of 3, and Slim Fly have almost no interference at two hops. Thus, one does not have to pay much attention to how to schedule pairs of communicating jobs, assuming they are executed on servers attached to routers separated by the distance l=2 (or l=1). The analysis of the path diversity between all router pairs facilitates reasoning about the expected latency of a topology with non-shortest path routing, for different router pairs, and offers hints on how to route. The majority of router pairs are connected by a number of different edge-disjoint paths that is larger than 1/3 of all their available network ports (indicated with circles and triangles). Such pairs do not require a lot of attention when selecting a path to route. The analysis of path interferences for shortest & non-shortest paths acilitates reasoning about the performance of both minimal and non minimal routing, and of scheduling of potential processes to different servers is presented in the last subfigure as 2D histograms.
5.1. Diversity of Shortest Paths
First, we consider the lengths and multiplicities of shortest paths. Understanding the former facilitates reasoning about the expected latency of a topology with shortest-path minimum routing, for different router pairs. Knowing the latter, on the other hand, enables taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing (Besta et al., 2020, 2021).
Definition 5.4.
For $s,t\in V$, the length of the shortest path $l_{\text{min}}(s,t)$ between them is
$l_{\text{min}}(s,t)=\min\left\{i\in\mathbb{N}:t\in h^{i}(\{s\})\right\}.$
Definition 5.5.
The shortest path multiplicity between $s,t\in V$ counts the number of shortest paths between $s$ and $t$ and is defined as
$n_{\text{min}}(s,t)=n_{l}(s,t)\text{ with }l=l_{\text{min}}(s,t).$
We provide example results in Figure 4, for both lengths of shortest paths (the left panel) and for their multiplicities (the right panel), for different sizes of four selected networks (indicated with different colors). The left panel illustrates the rich information about lengths of shortest paths. For example, most pairs are connected with a shortest path having two hops. Specifically, more router pairs in SF are connected with a 2-hop shortest path, than in HX2, for all topology sizes. This motivates more focus on making best-possible routing decisions over 2-hop paths in SF; 1-hop paths are negligible (especially for large networks). The right plot shows that, for most router pairs, the multiplicity of DF, HX2, and SF is low ($n_{\text{min}}\leq 2$) compared to HX3. We also observe that nearly all router pairs in SF have only 1 shortest path between them, making non-minimal multipathing crucial. Still, note that DF and SF host some outlier router pairs connected with large counts of shortest paths. Finally, most router pairs in HX have many shortest paths between them, and this ratio is growing with $N$, making ECMP a good option.
5.2. Diversity of Non-Shortest Paths
Next, we are interested in the number of edge disjoint paths between two vertices $s,t\in V$, that are not necessarily shortest. It facilitates reasoning about the expected latency of a topology with non-shortest path routing, for different router pairs, and offers hints on routing design.
For convenience, we define this measure for two groups of vertices $A,B$:
Definition 5.6.
The count of edge disjoint paths $c_{l}(A,B)$ at length $l$ between sets of routers $A\subset V$, $B\subset V$ is the smallest number of edges that can be removed such that $A$ and $B$ are not connected by any path of length at most $l$.
To obtain this count for two specific routers, we set $A=\{s\},B=\{t\}$. The above definition lets us easily derive the number of shortest edge disjoint paths between two vertices:
Definition 5.7.
The count of shortest edge disjoint paths $c_{\text{min}}(s,t)$ between $s,t\in V$ is defined as
$c_{\text{min}}(s,t)=c_{l}(\{s\},\{t\})\text{ with }l=l_{\text{min}}(s,t)$
$c_{\text{min}}(s,t)$, in contrast to $l_{\text{min}}(s,t)$, does not count shortest paths that share any edges. Hence, $c_{\text{min}}(s,t)\leq l_{\text{min}}(s,t)$.
We show example results in Figure 5. The top-left panel contains the histograms of $c_{l}$ at various lengths for DF, SF, and two HX instances. The values of $c_{l}$ of DF are in general lower than in the low dimensional HX instances (HX2 and HX3) and SF. Accordingly, $c_{l}$ of the HX and SF instances attains the network radix at lower lengths. HX2 has very high path diversity that is also “stable” (20 for all router pairs) for all higher path lengths ($l=3,4,5$). This facilitates the design for complex adaptive routing that routes around hotspots, uses complex paths, et cetera. Next, DF has relatively large diversity of paths for $l=5$, suggesting similar insights as for HX2, but the count is 11, so less than 20 by nearly 50%. Then, for $l=4$ or $l=3$, there is quite some diversity but the counts for different pairs may differ significantly, requiring attention when developing routing.
We further investigate $c_{l}$ with a detailed view of the count of edge disjoint paths for every router pair in a small SF, showing pairs with $\frac{c_{l}}{k^{\prime}}\leq 75\%$ (the bottom-left panel). This plot reveals all pairs of routers that do not come with a large path diversity between them. This enables a fine-grained performance debugging of routing and job scheduling schemes. The results show that the majority of router pairs are connected by a number of different edge-disjoint paths that is larger than 1/3 of all their available network ports (indicated with dark shade). Such pairs do not require a lot of attention when selecting a path to route. However, certain pairs do not come with large path diversity (cf. light shade). These specific pairs are not good candidates for scheduling communication-intense jobs.
5.3. Path Interference
So far, we considered measures that only take into account paths between two vertices. To capture the interactions of paths (of different lengths) between various pairs of vertices, we now define path interference. Here, paths between two router pairs $a,b$ and $c,d$ interfere if their total count $c_{l}(\{a,c\},\{b,d\})$ of edge disjoint paths at length $l$ is lower than the sum of the single destination count of edge disjoint paths $c_{l}(\{a,c\},\{b\})+c_{l}(\{a,c\},\{d\})$.
Definition 5.8.
Path interference $I^{l}_{ac,bd}$ between two router pairs $a,b$ and $c,d$ is defined as
$I^{l}_{ac,bd}=c_{l}(\{a,c\},\{b\})+c_{l}(\{a,c\},\{d\})-c_{l}(\{a,c\},\{b,d\})$
Understanding path interference facilitates reasoning about the performance of both minimal and non-minimal routing, and the scheduling of potential processes to different servers.
Note that $I^{l}_{ac,bd}\geq 0$, where equality is achieved when the disjoint paths from $a,c$ to $b$ do not share any edges with the disjoint paths from $a,c$ to $d$. Thus, if $I^{l}_{ac,bd}=0$, then connections can use fully the available link bandwidth. On the other hand, the higher the value of $I^{l}_{ac,bd}$, the more edges overlap so that the connections have to share this bandwidth.
Figure 5 (the top-right panel) shows histograms of interference $I^{l}_{ac,bd}$ up to length $l=5$ for SF, DF, and two instances of HX. At length $l=4$ and $l=5$ DF shows higher interference than HX2, HX3 and SF in contrast to length $l=3$ where interference of DF is lower. This insight could be used when deciding on non-minimal adaptive routing decisions: taking paths of length 3 is more beneficial than 4 or 5. Also, DF, HX3, and SF have almost no interference at two hops ($l=2$). Thus, one does not have to pay much attention to how to schedule pairs of communicating jobs, assuming they are executed on servers attached to routers separated by the distance $l=2$ (or $l=1$). Moreover, DF has also little interference between router pairs separated by 3 hops. Finally, at $l=3$ (i.e., for 3 inter-router hops), HX2, HX3, and SF have significant amounts of path interference, calling for care if devising routing/scheduling with such path lengths.
We further show a more detailed view of interference up to length $l=5$ on the bottom-right panel. Each single histogram illustrates the counts of router pairs experiencing a given scope of interference, i.e., their actual path interference versus the potential full (non-interfering) count of edge-disjoint paths, for a given length (as determined by the column) and for a given network (as determined by the row). Here, our pipeline also enables delivering various performance related insights. In general, we observe that the further a data point is from the “diagonal” of each histogram, the more path interference this data point indicates. Such far data points denote a large difference between the ideal amount of path diversity (on the X axis) versus the actual amount of diversity (on the Y axis). No topology is much imbalanced in this respect, as all data points far from the diagonal are dark (indicating few router pairs characterized a given difference between ideal and actual path diversity). Such pairs still require care when scheduling communication-intense jobs.
5.4. Computing Path Diversity Efficiently
In EvalNet, we primarily use algebraic algorithms to compute path diversity metrics, as they offer better scalability and parallelizability compared to purely combinatorial approaches, particularly for large, low-diameter networks (Besta et al., 2017; Solomonik et al., 2017; Solomonik and Demmel, 2011; Solomonik et al., 2014). For example, for counting the number of paths of fixed length between all node pairs, we use adjacency matrix exponentiation, where the $\ell$-th power of the adjacency matrix $A$ gives the number of $-\ell$ paths via $(A^{\ell}){i,j}$. This computation is efficient for small diameters and can be adapted for next-hop tracking using semiring-based variants. For length-restricted connectivity, we extend the algebraic path propagation algorithm of Cheung et al. (Cheung et al., 2011). For other metrics, we also explore max-flow-based approaches such as Push-Relabel and Ford-Fulkerson with BFS path constraints (used for sampled length-limited disjoint path analysis), as well as classic Gomory-Hu trees for all-pairs edge connectivity. While these combinatorial algorithms serve as correctness baselines or targeted tools for small-scale samples, EvalNet’s core design centers on the algebraic methods due to their efficiency, extensibility, and compatibility with parallel computation.
Figure 6. (left) Heatmap showing the Maximum Achievable Throughout (MAT), obtained with TopoBench, for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern. (right) Packet-based simulation runtimes [cycles] with Booksim2, for networks of 50k endpoints, for two injection rates, 0.01 and 0.5. As expected, networks with lower diameter usually provide lower latency (load=0.01). For load=0.5, the network global bandwidth starts to play an important role; here, DF has higher latency than FT3 because it has much lower global bandwidth.
[…]
[Figure is fully described in the text]Figure is fully described in the text.
6. FLOW-LEVEL SIMULATIONS
Flow-level simulators enable evaluation of large-scale systems, at a cost of a coarse design that may hinder realistic insights into performance. Flow-level simulations assume a continuous, real-valued flow model. Each flow has a rate $\mu\in\mathbb{R^{+}}$. The sum of all flows on each link must not exceed the link capacity. Without any additional constraints, this leads to constrained max-flow solutions in the topology graph. Other constrains (e.g., on the routes that can be taken by specific flows) lead to more realistic upper-bounds for performance of various routing schemes on specific topologies. However, since these models do not represent time accurately, they are unable to estimate latency or provide realistic insights into the impact of routing choices on flow control. As such, they are a useful tool for initial insights when designing a network.
For flow simulations, we provide an interface to TopoBench, an evaluation tool (Jyothi et al., 2016) that uses linear programming (LP) to derive MAT. Our TopoBench version supports state-of-the-art layered routing schemes such as FatPaths (Besta et al., 2020) or extended layered routing (Blach et al., 2024), built specifically for enabling non-minimal multi-path routing. Example results are in the left part of Figure 6, where we analyze a variant of FatPaths that offers non-minimal multipath routing. That part contains a heatmap showing MAT for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern. “#R” (Y axis) as well as “#P” (X axis) are parameters of the FatPaths routing scheme; they can be tuned for a specific topology. Usually, the higher #R is, the better the throughput becomes (because more layers entail more potential for multipathing (Besta et al., 2020)). Simultaneously, higher #R indicates more HW costs because one needs larger forwarding tables. Then, higher #P does not always enhance MAT (as confirmed by the results), because certain value ranges may cause larger path interference across routing layers, ultimately diminishing performance (Besta et al., 2020). Overall, for certain configurations, FatPaths on SF achieves very high MAT.
7. PACKET-LEVEL SIMULATIONS
Packet-level simulators offer a more detailed packet-based model of the network and routing. All actions in the simulation are modeled as events, which are scheduled to happen at a given point in time. For each event, some code is executed, which can update the state of the simulation, and produce new events that are scheduled to happen in the future. The main simulation loop keeps the scheduled events in a priority queue and advances the simulation by dequeuing and executing successive events.
7.1. Which Simulator to Use?
A large number of simulators exist, and it is challenging to select the one best suited for one’s purposes. We analyze a broad selection of available simulators, and show that none scale to the desired sizes of hundreds of thousands of servers. The overview is in Table 2. We analyzed their codebases, investigating which (1) facilitate modifications that could enhance scalability, (2) are easy to use and extend without requiring distributed-memory infrastructure support, and (3) cover a large number of protocols and settings. Based on that analysis, we selected two simulators: htsim (Raiciu et al., 2010) for lossy TCP/IP stacks and Booksim2 (Jiang et al., 2013) for lossless HPC architectures. Other simulators do not meet these criteria. For example, SST and ROSS natively target HPC clusters and we concluded that they are relatively complex to set up, use, and extend on low-cost PCs.
htsim The htsim simulator (Raiciu et al., 2010) provides a lightweight infrastructure for lossy networks. It only models the transport layer: there is no model for links or routers. Instead, the route for each packet is pre-computed as a sequence of queues that the packet will pass through. Such a route specification is attached to the packet. Each queue has a finite service rate, which models the link capacity. Due to this design, htsim is highly flexible, and new topologies can be added straightforwardly: the topology only affects the route computation, which is explicitly called for every flow during its setup. The disadvantage is that many adaptive routing schemes cannot be modeled, since there is no per-router state that could affect the routing. Furthermore, all flows are initialized before the first event is processed, and all routes are kept in memory. EvalNet offers a seamless interface to htsim. We select htsim as it targets TCP stacks; it supports by default modern TCP-based schemes such as NDP (Handley et al., 2017), and it can be easily extended with new routing schemes.
Booksim2 Booksim2 targets lossless networks. It explicitly models input-queued routers, thus one can easily extend the simulator with new routing protocols. Moreover, there is support for flow control unit (flit) packets, and thus for schemes such as wormhole routing or virtual cut-through flow control. Virtual channels are also supported. One can specify the total buffering/port, router delay for credit processing, delays for channel latency, router and VC allocations, and processing in a crossbar. As with htsim, EvalNet offers modules that enable seamless evaluation of any generated topologies within the Booksim2 setting.
Simulator Scalability [#servers] Design
LogGOPSim (Hoefler et al., 2017) 1024 (Hoefler et al., 2017) SM
htsim (Raiciu et al., 2010) 128 (Raiciu et al., 2010) SM
NS2 (Issariyakul and Hossain, 2009) $<$1,000 SM
booksim (Jiang et al., 2013) 10,000 (Besta and Hoefler, 2014) SM
FOGSim (Garcıa et al., 2014) 16,512 (Benito et al., 2015) SM
OMNet++(Varga and Hornig, 2008) 100,000 (Gamer and Mayer, 2009) SM
NS3 (of Washington NS-3 Consortium, 2024) 5,000 (Pelkey and Riley, 2011) DM
NS4 (Fan et al., 2017) N.A. DM
BigSim (Zheng et al., 2004) 46,656 (Jain et al., 2016) DM
SST (Rodrigues et al., 2016) 110,592 (Groves et al., 2016) DM
xSim (Böhm and Engelmann, 2011) 2,097,152 (Engelmann, 2014) DM
ROSS/CODES (Cope et al., 2011) 1,000,000 (Wolfe et al., 2016) (Slim Fly), 50,000,000 (Mubarak et al., 2012) (Dragonfly) DM
htsim [this work] 1,000,000 SM
Booksim2 [this work] 50,000 SM
Table 2. The comparison of available packet-level network simulators. “Scalability” is the largest size of a topology (simulated with a specific simulator) that we were able to find in the literature. “Design” includes the details of the compute platform used for simulations in the “Scalability” column: “SM” (shared-memory design), “DM” (distributed-memory design).
Element Range (lower–upper bounds) 10,000 servers 100,000 servers 1,000,000 servers
Count (relative) Count (total) Memory (per item) Count (relative) Count (total) Memory (total) Count (relative) Count (total) Memory (total) Count (relative) Count (total) Memory (total)
Network router 1 $10^{2}$$10^{4}$ (1) 1 242 (1) 1 1,058 (1) 1 5,618 (1)
cable (per router) 10–1000 $10^{4}$$10^{6}$ (1) (4) 8.5+40 11,737 (1) (4) 17.5+90 113,735 (1) (4) 39.5+200 1,345,511 (1)
server (per router) 10–1000 $10^{4}$$10^{6}$ 5kB 40 9,680 47MB 90 95,220 460MB 200 1,123,600 5.4GB
routing entries $O\left(N_{r}^{2}\right)$ $10^{4}$$10^{8}$ (2) 100B $O\left(N_{r}^{2}\right)$ 58,564 6MB $O\left(N_{r}^{2}\right)$ 1,119,364 110MB $O\left(N_{r}^{2}\right)$ 31,561,924 3GB
Workload task (per server) 1-10 $10^{2}$$10^{5}$ (1) 1 968,0 (1) 1 95,220 (1) 1 1,123,600 (1)
flow (per task) 1-100 $10^{2}$$10^{7}$ 2kB 100 968,000 1.8GB 10 952,200 1.8GB 1 1,123,600 2.1GB
path (per flow) 1-10 $10^{2}$$10^{8}$ 600B 5 4,840,000 2.7GB 5 4,761,000 2.7MB 5 5,618,000 3.2GB
packet (per flow) 10-1000 $10^{3}$$10^{10}$ (3) (5) 110 106,480,000 (3) (5) 110 104,742,000 (3) (5) 110 123,596,000 (3)
Table 3. Estimation of the number of elements of a data center Slim Fly simulation. The numbers and thus required resources vary widely depending on a chosen configuration. “Range” lists parameter bounds considered in this work, in the following columns we show numbers for example configurations. Memory estimates are based on htsim. The chosen workload parameters keep the simulations feasible on a laptop (configuration details are in 7.3). 1 Negligible for relevant configurations. 2 Data depends on the routing scheme. 3 Objects are transient and accounted for in the flow counts. 4 Inter-router cables + server links. 5 The average over the flow size distribution, excluding re-transmissions.
7.2. Can We Simulate Large Networks on Low-Cost HW?
Next, we conduct a feasibility analysis in which we argue that million-server packet-level simulators should in theory be achievable on a low-cost PC. For this, we analyze the approximate memory and time cost of targeted large-scale simulations.
Number of Simulation Elements We estimate the number of elements of a simulated topology (e.g., servers) and of a simulated workload (e.g., flows which require flow control state). The results of the analysis are in Table 3. The number of network elements does not pose the most serious scalability problems. Instead, the number and corresponding size of elements related to workloads dominates the memory usage and the running time.
Memory and Time Requirements The total memory usage and simulation time depend on the simulation software and on granularity, i.e., what elements of the simulated workload are explicitly stored and simulated. We estimate these numbers in Table 3, which also shows what may not be feasible in the largest-scale simulations. Specifically, we cannot simulate large workloads on large networks, and must not store any per-packet state.
The lowest required amount of memory is determined by the number of simulated flows, because we want to simulate end-to-end flow control for each flow. Thus, the flow control algorithm state for each flow needs to be stored. We also account for the packets in flight on a per-flow basis, since the memory occupied by packets is tightly coupled to the number of concurrent flows. Memory consumed is about 2kB/flow plus 600B/path, which could probably be further reduced. To make large simulations feasible, it is crucial to keep the per-flow overhead low. Special attention needs to be paid to monitoring and debugging code, which can easily dominate the per-flow state. For very large networks, the quadratically increasing size of the routing tables starts to dominate, but as Table 3 shows, this was not a problem for our simulations.
The lowest amount of time is determined by the amount of packet forwarding. Each packet forwarding has to be processed as at least one event, and these events constitute the majority of simulation time. In the modern low-diameter networks, each packet has to be forwarded $\approx$2–4 times to reach its destination, and we observed around 60 events processed for each transmitted packet (this includes entering and leaving queues on the path through the network, for the data packet as well as the ACK packet and potential retransmissions). We observe event rates of around $10^{6}$ per second on one CPU core. Since each event likely causes at least one cache miss, we cannot expect much higher rates without distributed simulation or middle- to high-cost PC stations, which are outside the scope of this work.
7.3. How to Simulate Really Large Networks on Low-Cost HW?
Finally, we show how to use htsim (Raiciu et al., 2010) and Booksim2 (Jiang et al., 2013) to enable more scalability. In Booksim2, the key limiting factor that we identified was processing related to single packets. This processing is usually related to gathering statistics, which is not relevant for the actual simulation outcomes. In Booksim2, eliminating this processing was the main enabler for being able to simulate networks with 10,000 endpoints and more. In htsim, on the contrary, the main bottleneck was the storage related to flows. Specifically, we had to add a more efficient, routing table based scheme for arbitrary topologies. We do not use the provided logging solution, which tends to produce too much output at the simulation scales that we consider. Instead, the statistics of each flow are printed in text format to standard output when the flow finishes. Another scalability obstacle was the net_paths structure present in the sample programs, a preallocated $N^{2}$ size matrix of routes, which would only be sparsely populated for permutation workloads and would dominate memory use for large networks.
7.4. Representative Simulation Results
We simulate networks with $\approx$10k, $\approx$100k, and $\approx$1M servers using a laptop with 16GB RAM and an Intel Core i7-8550U CPU; each run takes 1–3 hours. Networks are 5$\times$ oversubscribed, as in configurations studied in (Kassing et al., 2017), and include topologies larger than those used in prior large-scale SF simulations. Due to the cost of simulating large networks, we capture only a few milliseconds of activity—sufficient for reliable relative performance comparisons. We focus on flow completion time (FCT) distributions as a function of flow size, using the pFabric web search distribution (average $\approx$1MB) and injection rates of 40–60 flows/server/s. Results are shown as mean, 10th, and 99th percentiles. To mitigate artifacts from the fixed injection window, we carefully interpret short and long flows.
Figure 7. Results of large-scale simulations.
[…]
[Comparison of the Jellyfish and Slim Fly topologies]The figure shows three different plots comparing various aspects of the Jellyfish and Slim Fly topologies. The first plot presents the flow completion time with similar parameters as in the previous figure: Jellyfish and Slim are identical for the 10% percentile and the mean, but Jellyfish has a smaller completion time for the 99% percentile. The second subfigure plots the distribution as a histogram for 10.000, 100.000 and 1.000.000 servers for long flows (200KB) and an arrival rate of 50/s. Jellyfish exhibits a better tail latency. The third plot compares the flow completion time with respect to the arrival rate (ranging from 40 to 60) of long flows and the number of servers (10.000, 100.000 and 1.000.000). Both topologies exhibit the same mean performance, but Jellyfish is faster for the 10% and 90% percentile.
As an example, we analyze 10k, 100k, and 1M server simulations, see Figure 7. The left plot shows FCT as a function of size. For long flows ($v>200$kB), FCT is limited by bandwidth, while for shorter flows, latency dominates. The middle plot illustrates the distribution of FCTs for middle-size flows: here, the better tail behavior of the randomized Jellyfish topology is clearly visible. In the right plot, we show the impact of the flow arrival rate $\lambda$. Even a small increase in load causes a large degradation in FCT in the considered networks, due to the used oversubscription. The 1M server simulation is less affected due to the limited warmup time.
Our analysis illustrates that million-server packet-level simulations on a simple low-cost laptop are feasible. However, they require some compromises on the quality of simulation, most importantly on the number of allowed flows and on the methodology (i.e., simulating a fixed amount of flows instead of a fixed time window after reaching a steady-state).
8. COMPREHENSIVE ANALYSIS & TAKEAWAYS
Topology Lat. GlB. Ext. C&P MS MnS
[…]
EvalNet enables multifaceted analysis of numerous network topologies. We list the considered networks in Table 4, and provide an assessment of their important properties (performance, cost & power consumption, path diversity), obtained using EvalNet. The table compiles the results described in the previous sections of this paper, as well as further results illustrated in the Appendix in Figures 856.
8.1. General Observations & Patterns
A central observation is the critical role of path diversity (both shortest-path (MS) and non-shortest-path (MnS)) in determining a network’s performance and scalability. Topologies such as Xpander and OFT exhibit high MS, meaning they provide many equal-cost paths between node pairs. This allows routing schemes like ECMP to distribute traffic efficiently, reducing congestion and delivering both low latency and high global bandwidth under minimal routing. These networks are thus well-suited for environments that rely on deterministic routing, where fault tolerance and load balancing are achieved through structural redundancy alone.
In contrast, Moore-bound diameter-2 networks like SF and PF attain excellent performance through topological minimalism (i.e., by minimizing average path length and diameter). However, their MS is low: most router pairs have only a single shortest path, rendering traditional ECMP ineffective. These networks compensate by offering very high MnS, which enables advanced non-minimal multipathing strategies such as Valiant load balancing and FatPaths, allowing traffic to be dispersed over longer, disjoint paths to avoid congestion. This tradeoff is critical: although SF, PF and other such networks lack diversity in minimal paths, their rich non-minimal structure, if properly exploited, enables them to outperform MS-rich networks under high loads—provided the routing stack supports sufficient adaptivity and complexity (e.g., congestion-aware routing, larger tables, multipath tracking).
This divergence between MS and MnS has a broader design tradeoff. Networks with high MS (e.g., Xpander) typically have slightly larger diameters but benefit from simplicity and robustness of minimal routing. Networks with low MS but high MnS (e.g., SF, PF) require intelligent routing to achieve high performance but can then perform remarkably well. On the other hand, low-dimensional regular topologies such as torus and mesh suffer from low values of both MS and MnS. While these networks are easily extensible in terms of size, their limited path diversity leads to high latency. Without structural augmentation (e.g., express links or randomized shortcuts), their performance does not scale with node count.
Tree-based networks, including Fat Trees, k-ary n-trees, and eXtended Generalized Fat Trees, offer a moderate level of both MS and MnS due to the presence of multiple upper-layer routes. These designs are non-blocking when fully provisioned and can deliver high global bandwidth by supporting ECMP and basic multipathing. However, they suffer from moderate-to-high latency due to their hierarchical depth, and may lack resilience of the more modern high-MS or high-MnS topologies that have good expansion properties.
These insights tie into performance, cost, and scalability tradeoffs. While SF and PF achieve top-tier latency and bandwidth, their fixed diameter fundamentally limits extensibility. Namely, they cannot grow beyond a certain size without losing their optimal structural properties. In contrast, torus and mesh networks offer near-unlimited extensibility due to their non-constant diameter, but this comes at the cost of poor latency and reduced throughput. Topologies with very high radix, like HyperX and Multi-Layer Full Mesh, provide rich connectivity and sometimes high MS/MnS, but incur substantially higher construction and power costs, as they require more expensive switching hardware due to very high radix. Dragonfly and Cascade Dragonfly offer a middle ground, balancing latency, cost, and path diversity, which explains their popularity in practical deployments.
Ultimately, while our findings highlight that each topology occupies a different point in the design tradeoff space, low-diameter networks such as Slim Fly and PolarFly emerge as particularly desirable candidates. Their minimal path lengths enable very low latency and bandwidth, which are critical for modern high-performance workloads. Although they offer limited shortest-path diversity, their rich structure of non-minimal paths allows them to sustain high throughput under load—provided that routing schemes can exploit this flexibility. EvalNet enables these tradeoffs to be quantified and visualized precisely, supporting informed design decisions based on latency, throughput, scalability, cost-efficiency, and path diversity.
8.2. Example Detailed Insights
We also provide a few selected more detailed insights that could be used, e.g., as a basis for enhancements of routing strategies. Full results and observations are in the Appendix.
Topologies based on algebraic expanders minimize contention for multipathing. For example, Spectralfly’s path interference is nearly zero, even matching random Jellyfish variants, implying that one can exploit non‐minimal detours almost for free, boosting available parallel paths without contention.
Express Meshes have bimodal latency. Express Meshes have a “gap” in their shortest-path multiplicity: most node-pairs have very few ($\leq$2) or a large number ($\geq$5) of disjoint shortest paths, with almost no pairs in between. Hence, under uniform ECMP, about half the traffic has rich bandwidth, while the other half is starved, creating a bimodal latency distribution (unless one employs adaptivity).
Megafly has hop-bound tail-latency. Megafly variants have little path interference for up to 3 hops, yet develop a small set ($\approx$10) of high-interference outliers at more than 3 hops. Hence, one can exploit non-minimal multipathing up to 3 hops with near-ideal bandwidth, but more hops risk large tail-latency spikes.
PolarStar variants have performance trade-offs. PolarStars with Paley-subgraphs have marginally higher shortest-path multiplicity at 5 hops and only moderate path interference, whereas PolarStars with BDF-subgraphs yield lower overall interference (smoother latency) but slightly fewer total paths. Thus, one can tune PolarStar for either peak throughput (Paley) or tighter latency tails (BDF) based on application’s burstiness.
9. CONCLUSION
In this work, we introduce EvalNet, a comprehensive toolchain for the generation and in-depth analysis of extreme-scale interconnection networks, with a strong focus on path diversity as a core enabler of performance and scalability. EvalNet integrates topology generation, fine-grained path diversity analysis, performance modeling, and cost evaluation into a unified and extensible framework. Using EvalNet, we conduct the widest comparative study to date across 25 network topologies, revealing how shortest- and non-shortest-path diversity relate to key performance and deployment metrics such as latency, throughput, cost, and power. Our analysis shows that the ability to exploit non-minimal multipathing is essential for maintaining high throughput in low-diameter, scalable networks. EvalNet interfaces seamlessly with established simulators and serves as a practical and versatile tool for guiding future innovations in large-scale network design in data centers and supercomputers.
Acknowledgments
We thank Haoran Zhao for his code contributions. We thank Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). This project was supported by JSPS KAKENHI Grant Number JP19H04119.
Appendix A Appendix
Figure 8. Shortest path plot for Kautz and Arrangement graphs. Kautz graphs have lower shortest paths compared to their Jellyfish equivalent for small networks. They also have lower shortest paths when compared to a small Arrangement network. A small Kautz graph has only a diameter of 3, while the Arrangement graph has already a diameter $\geq$ 5. The analysis of the Arrangement graphs compared to their Jellyfish equilibrium, which is a Jellyfish topology built with the same hardware, looks fairly similar. The Jellyfish equilibrium has less shortest paths of higher lengths $l$ = 5. The amount of shortest paths of low length (between 1 and 3) gets smaller, the bigger the network gets for all topologies. For Kautz graphs, these drop-offs are much more pronounced. There are about 100 times fewer shortest paths of lengths 1 when we compare a small Kautz graph with a large Kautz graph. In comparison, Arrangement networks have a drop-off between 5 and 10 times.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 9. Multiplicity plot for Kautz and Arrangement graphs. For Arrangement networks, the shortest path density first decreases, then increases. In contrast, Kautz graphs have about the same amount of path multiplicity, no matter the value of $n_{min}$. In the large Kautz network, we can see a significant drop off in the amount of multiplicity. This drop off is less significant for $K_{6,n}$. For Arrangement networks, these numbers of multiplicities stay about the same. There is only a small shift, for $n_{min}$ = 1 the number of shortest path multiplicities sinks for higher N, but for $n_{min}$ = 4 or 5 they only sink slightly. Kautz graphs have a gap in the multiplicity counts. They either have a low amount (1 or 2) of shortest paths between any two nodes or a lot of overlap ($\geq$ 4) in the multiplicity. The analysis of the Jellyfish equivalent looks similar to the Arrangement networks. The Jellyfish equivalent of the Kautz graphs on the other hand do not have this gap in the multiplicity. For $K_{4,n}$ and $K_{5,n}$ the drop off in the multiplicity count can also be seen in the Jellyfish equivalent for N $\approx$ 100,000. This cannot be seen for the Jellyfish equilibrium for $K_{6,n}$.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 10. Shortest path plot for mesh topology of different dimensions. The number of shortest paths increases while $l_{min}$ increases. For bigger networks, there are fewer shortest paths with a small value. The Jellyfish equivalent analysis looks fairly similar compared to the meshes. Only for $l_{min}$ = 5, the Jellyfish network with higher N does not strictly have fewer amounts of shortest paths.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 11. Multiplicity plot for meshes of different dimensions. As expected, the analysis looks fairly similar across all dimensions. For all sizes and all dimensions, the multiplicity count is the highest for $n_{min}\geq$ 5. For bigger network sizes, the amount of shortest path multiplicity decreases with roughly the same factor, independent of $n_{min}$. In contrast, the Jellyfish equivalent networks have almost no difference in the multiplicity count for $n_{min}$ = 1. For bigger values of $n_{min}$, the shortest path multiplicity drops even more compared to the original mesh, when we increase the network size.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 12. Shortest path plot for the 4D Express Meshes with different gap sizes. We choose the lowest g, such that for each Express Mesh 4DeMESHx we have at least x additional express connections per dimension. As one can expect, the Express Meshes have lower shortest paths than the original mesh due to their express connections. For some network sizes, the Express Meshes have only a diameter of 4, while the standard 4D mesh has a diameter of $\geq$ 5. While the number of shortest paths increase for bigger $l$ in the Express Meshes, their equivalent Jellyfishes don’t show a similar pattern. They increase first before they decrease for the biggest value of $l$ where there still exists a shortest path.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 13. Multiplicity plot for Express Meshes of dimension 4. For small and medium-sized networks, there is a gap in the multiplicity counts. The length of the shortest path for any two nodes exists either once or twice, or $\geq$ 5 times. For large Express Meshes, this gap is not visible anymore. Express Meshes keep the high multiplicity of shortest paths, even for higher N, due to their express connections, especially for lower gap sizes. For all network sizes, the Express Meshes have the highest shortest paths multiplicity count for $n_{min}\geq$ 5. There is a big difference in the multiplicity of the meshes compared to the equivalent Jellyfish. In contrast to the base Express Meshes, the Jellyfish equilibriums are well distributed along all values for $n_{min}$ for small and medium networks.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 14. Shortest path plot for 6D Express Meshes with different gap sizes. We chose the lowest g, such that for each Express Mesh 6DeMESHx we have at least x additional express connections per dimension. The shortest path analysis of these Express Meshes look similar to the analysis of their 4D counterpart. Express connections do not allow the 6D graph to become a diameter $\leq$ 5 network, since any path only traverses over one dimension. The Jellyfish equilibrium is also fairly similar when comparing the 6D to the 4D analysis. The amount of shortest paths grows as $l$ grows, before dropping down for some outliers that have a shortest path length of 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 15. Multiplicity plot for Express Meshes of dimension 6. We can see again the gap in the amount of shortest path multiplicity. There either exist very few paths of length equal to the shortest path between two nodes, or there exists a lot of them (more than 4). Only the large Express Mesh with two additional express connections per dimension does not show this gap. Same as in the 4D case, most shortest paths have a multiplicity $\geq$ 5. In case of the Jellyfish equilibrium, the shortest path multiplicity count drops at higher $n_{min}$. Just for $n_{min}$ = 5, the multiplicity count increases, as they have as well the most shortest paths with multiplicity $\geq$ 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 16. Shortest path plot for the Tofu topology, 6D torus and 6D mesh. As we have already seen in the analysis of the meshes, the amount of shortest paths sinks as the network size gets bigger. The bigger $l$ gets, the more shortest paths exist of length $l$. All networks show similar patterns. Tofu has more shortest paths $l_{min}\leq$ 5 compared to all the other meshes for small networks, but is worse for bigger networks. The same is true when we compare Tofu to the equivalent Jellyfish.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 17. Multiplicity plot for the Tofu network, 6D torus and 6D mesh. Similar as for the meshes, all the topologies have the most multiplicity count where $n_{min}\geq$ 5. The Tofu network has even the highest multiplicity count for $n_{min}\geq$ 5 for all small networks. In medium and large networks, Tofu does not have the most multiplicity count anymore. The bigger the network sizes are, the smaller the multiplicity counts get. In the analysis of the torus, we can find the already mentioned gap in the shortest path multiplicities for small and medium-size networks. Tofu has this gap only for small networks. The Jellyfish equilibrium of the Tofu network is well distributed along all values for $n_{min}$ for small and medium networks. For the large network, there is a significant drop-off and the multiplicity is about 30 times smaller for $n_{min}\geq$ 5 compared to the small or medium network.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 18. Shortest path plot for PolarStar compared to Dragonfly and Slim Fly. We can clearly see that all are diameter $\leq$ 3 topologies. All of the topologies have in general lower smallest paths compared to their equivalent Jellyfish graph. Slim Fly has only shortest paths of length 2. All the others look fairly similar, having fewer shortest paths with $l_{min}$ = 1 compared to $l_{min}$ = 2 or 3. The amount of shortest paths of length 1 and 2 decreases as the network size increases. The shortest paths of length 3 slightly increase though. There is almost no difference for the analysis between PolarStar with BDF or Paley subgraphs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 19. Multiplicity plot for PolarStar, Dragonfly and Slim Fly. Slim Fly has the least multiple shortest paths of all topologies. Only for medium and large networks, there are pairs of nodes with shortest path multiplicity bigger than 3. All topologies have the most shortest paths only once between any two nodes. PolarStar with BDF subgraphs has only a few shortest path multiplicity count for $n_{min}$ = 3 or 4, while PolarStar with the Paley subgraphs has a well-distributed count for all values of $n_{min}$.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 20. Shortest path plot for Cascade Dragonfly, PolarStar and Spectralfly. The amount of shortest paths of length 1 and 2 decreases, if the network gets bigger, because a single node in a group can only reach 21 nodes from another group with at most 2 jumps. Meanwhile, all 95 nodes in the same group can be reached with 2 gaps. The amount of node pairs that have a shortest path $\geq$ 3 stays almost the same. Spectralfly has the least amount of shortest paths of length 1 for all network sizes. The amount of shortest paths of low length stays the same or gets even better for higher-size networks. The Jellyfish equivalent of the Cascade Dragonfly has only a diameter of 3 or 4 (depending on the network size), while Cascade Dragonfly is a diameter 5 topology. Spectralfly’s Jellyfish equilibrium on the other hand looks similar to its base counterpart. Only for large network sizes, the Jellyfish graph shows no shortest paths of length 5, while the Spectralfly graph still has some.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 21. Multiplicity plot for the Cascade Dragonfly, PolarStar and Spectralfly networks. Most node pairs in the Cascade Dragonfly topology have multiple ($\geq$ 5) shortest paths. The multiplicity count either stays about the same or gets bigger, as the Cascade Dragonfly graph increases its size. The small Spectralfly network has only multiplicity counts $\leq$ 4. For large and medium networks, the Spectralfly topology has slightly more multiplicity count for $n_{min}$, however only by a small margin. For small and medium networks, the equivalent Jellyfish graph of the Cascade Dragonfly has as well the most shortest path multiplicity for $n_{min}\geq$ 5. The large Jellyfish has about the same multiplicity count over all values of $n_{min}$. While the equivalent Jellyfish of Spectralfly drops in multiplicity counts for higher values of $n_{min}$, for medium and large networks it increases.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 22. Shortest path plot for Megafly with different values for g and Dragonfly. We can see, that Megafly is a diameter 3 network as well. Megafly has no shortest path of length 1, since every leaf router is only connected to spine routers, and not to another leaf router. Because two nodes have only shortest path of length $l$ = 2 if they are in the same group, most shortest paths are $l_{min}$ = 3 for Megafly. All topologies have the same amount of shortest paths of length 3 for all network sizes. The analysis of the equivalent Jellyfishes looks pretty similar compared to the base Megafly network, where Megafly has shortest paths of length $l$. The Jellyfish graphs have additional shortest paths of length 1 and sometimes length 4.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 23. Multiplicity plot for the Dragonfly and Megafly topologies with different values for g. All topologies have mostly only one shortest path between any two nodes. Between any two leaf nodes in Megafly, there is mostly only one shortest path. There is either one shortest path between the nodes or a bunch ($\geq$ 5) of them. Dragonfly has some shortest paths of different multiplicity, but most lie in the extremes $n_{min}$ = 1 or $n_{min}\geq$ 5. Their equivalent Jellyfish on the other hand have well-distributed shortest path multiplicities.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 24. Shortest path plot for Fat Trees and the k-ary 3-tree. All topologies exhibit the same patterns for the shortest paths of a given length. While the shortest paths of length 4 stay the same for any network size, the number of paths of length 2 decreases as the network gets bigger.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 25. Multiplicity plot for Fat Trees and the k-ary 3-tree. All networks and all sizes have a path multiplicity count of 5 or higher.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 26. Shortest path plot for length $l_{min}$ up to $l$ = 5 for multiple variants of xGFTs. There are only shortest paths of even length. There exist more shortest paths of length 4 than length 2 in general. As the network size increases, the number of shortest paths at each length decreases.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 27. Multiplicity plot for multiple variants of xGFTs. Variant 4 has the lowest amount of multiplicity count. It only has a multiplicity count of 2 and 4. For large network sizes, the 8S variant has even more multiplicity than the smaller networks of that variant.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 28. Edge disjoint path plots for Kautz and Arrangement graphs of small sizes. The connectivity for $A_{n,4}$ is better than the connectivity for $A_{n,5}$. Both have in general fewer edge disjoint paths than the Kautz topologies. $K_{6,n}$ has the best connectivity for these topologies. Furthermore, the Kautz networks analysis looks similar to its equivalent Jellyfish.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 29. Edge disjoint path plots for meshes of different dimensions. For all lengths $l$, the meshes have mostly zero edge disjoint paths between any two nodes. Furthermore, we can see that there are almost no node pairs, where there are 2 paths of a given length $l$. In contrast, none of the Jellyfish equivalent graphs have zero edge disjoint paths for $l$ = 5. The 4D mesh has a slightly worse path diversity than the 5D and 6D meshes.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 30. Edge disjoint path plots for the Tofu topology, 6D torus and 6D Express Mesh. Similar to the diversity analysis of the meshes, all the topologies have zero path diversity for most node pairs of any given length. For both Tofu and the 6D torus network, their path diversity counts lie in three groups. They have diversity counts of 1, $\approx$ 4 and $\approx$ 10. This is even more extreme for the 6D torus network and cannot be seen for the 6D mesh graph. Same as before, the equivalent Jellyfish graphs have a well-distributed diversity count.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 31. Edge disjoint path plots for Cascade Dragonfly, PolarStar and Spectralfly topologies. Both Spectralfly and Cascade Dragonfly have mostly a low diversity count. Cascade Dragonfly has an additional peak with node pairs of a higher diversity count. Even for length $l$ = 5, Spectralfly has only node pairs with 4 edge disjoint paths.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 32. Edge disjoint path plots for Dragonfly, PolarStar and Slim Fly topologies. Slim Fly has the most path diversity by far. It has a diversity count $\geq$ 10 for most nodes, even for paths of length 3. The diversity analysis of PolarStar is pretty similar to Dragonfly for length 1 to 4. For $l$ = 5 we can see that Dragonfly has a higher diversity count by 2 for all node pairs. The PolarStar graph with Paley subgraphs has slightly higher edge disjoint paths for $l$ = 5. Both variants of PolarStar have the same or better diversity compared to their Jellyfish equivalent for all lengths.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 33. Edge disjoint path plots for Dragonfly and Megafly networks with different values of g. All Megafly networks have a low path diversity of either 0 or 1 for most node pairs and all lengths. Furthermore, all have a second peak for a very high diversity count. Depending on the value of g, this peak is between 7 and 15. Although Megafly has node pairs with a higher diversity count, Dragonfly has a well-distributed diversity count between 1 and 5. It is difficult to compare the values with this plot. The diversity counts of all Jellyfish equilibriums show a Gauss distribution for $l$ = 3.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 34. Edge disjoint path plots for the Fat Trees and k-ary 3-tree. All the topologies have only specific values for the path diversity. None of the topologies have any node pairs that have a diversity count of zero for $l\geq$ 4. The k-ary 3-tree has the best path diversity among all other topologies.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 35. Edge disjoint path plots for the xGFTs of different variants. There exist only edge disjoint paths of specific values. For $l\geq$ 4, xGFT8 does not have any edge pairs with a diversity count of zero anymore. The k-ary n-tree has edge nodes with a higher diversity count than any other topology, but it has still node pairs with a diversity count of zero for $l$ = 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 36. Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 3. All router pairs have a connectivity $\leq$ 75% for $l$ = 3. We can see the segments (blue) of length 4 connecting a node to its neighbours. This can be seen horizontally as well as vertically. Due to this mesh of edges, all nodes are connected and there exist only a couple of node pairs with connectivity $<$ 25%.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 37. Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 5. All router pairs have a connectivity $\geq$ 75% for $l$ = 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 38. Low connectivity plot for the PolarStar graph of length $l$ = 3. The structure graph is an ER2 graph and the subgraph is a BDF3 graph. We can see the 7 well-connected BDF3 clusters of each 8 nodes in the diagonal. The other well-connected areas exist due to the connection of the ER2 graph. Every node in 0 - 7 is connected to a node in 31 - 39, 40 - 47 and 48 - 55.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 39. Low connectivity plot for the PolarStar graph for the length $l$ = 5. The structure graph is an ER2 graph and the subgraph is a BDF3 graph. We can clearly see, that for $l$ = 5 this PolarStar graph has a good connectivity $\geq$ 75% for all router pairs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 40. Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 3. Each cluster has a good connectivity within itself. On the diagonal, we can also see the 16 chassis for the 5 levels. The connectivity between the two clusters is very low and only about half of the nodes in each cluster have a path to nodes in the other cluster.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 41. Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 5. Right underneath the first cluster, we can see the four node pairs that connect the two clusters. Even for paths of $l$ = 5, the connectivity between the two clusters stays low at only 10%.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 42. Low connectivity plot for the Tofu topology with X = Y = Z = 2 and $l$ = 3. Not all nodes are connected with each other at length $l$ = 3. The 8 clusters of 12 nodes each are visible in the diagonal. The small blue diagonal lines show the edges between the nodes in different clusters. Each row is intersected by 3 of these lines. This is the number of connections each node has to other clusters.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 43. Low connectivity plot for the Tofu topology with X = Y = Z = 2 and length $l$ = 6. There is a good connectivity $\geq$ 70% for most node pairs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 44. Interference plots for Kautz and Arrangement graphs of length $l$ = 5. The Arrangement networks have the same or less interference than their Jellyfish equivalent counterpart for $l\leq$ 5. For $l$ = 5, both have slightly more interference. The same can be seen for $K_{4,n}$ and $K_{5,n}$ Kautz graphs, but not for $K_{6,n}$, which has the most interference for $l$ = 3 and $l$ = 4, but the least interference for $l$ = 5.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 45. Two-dimensional detailed interference plots for Kautz and Arrangement graphs of length $l$ up to 5. $A_{n,4}$ has more interference than $A_{n,5}$. The different values of k for $K_{b,n}$ Kautz graphs have less of an impact on the amount of interference, but at which length $l$ they will occur. Most of these networks have the most interference at $l$ = 4. Their equivalent Jellyfish have similar interference analysis, but in general a bit less.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 46. Interference plots for meshes of different dimensions of lengths up to $l$ = 5. While 4D and 6D meshes have a pretty similar interference analysis, the 5D mesh has more interference. Furthermore, for $l\leq$ 4, the base networks of all the meshes have higher or equal interference than their equivalent Jellyfish. 4D and 6D meshes have better interference for $l$ = 5, which is not true for the 5D mesh. For $l$ = 5, all meshes have more node pairs with zero interference compared to their equivalent Jellyfish graphs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 47. Two-dimensional detailed interference plots for meshes of different dimensions of lengths up to $l$ = 5. We can see, that all meshes have similar interferences analysis. For the base mesh networks and $l$ = 5, all components are affected by interference, while for the equivalent Jellyfish, only a subset of them are.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 48. Interference plots for Tofu, 6D mesh and 6D torus with lengths up to $l$ = 5. The Tofu topology has about the same amount of interference as the 6D torus network, but more than the regular 6D mesh. For $l$ = 5, both torus and the Tofu graph have more interference than the 6D mesh and more than any of the Jellyfish equivalents.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 49. Interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l$ = 1 up to $l$ = 5. Slim Fly has the most interference at $l$ = 3, but less interference for all other values of $l$. The other topologies have the most interference for $l$ = 4. The interference analysis looks similar for the topologies when comparing them to their equivalent Jellyfish, although all of the base topologies have slightly more interference.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 50. Two-dimensional detailed interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l\leq$ 5. The detailed analysis of both PolarStar graphs are fairly similar. While for Dragonfly, the interference for $l$ = 4 is contributed by a lot of components, only a small amount of components contribute to the high amount of interference for PolarStar.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 51. Interference plots for Cascade Dragonfly, Spectralfly and PolarStar with both subgraphs of length $l$ = 5. Spectralfly has little interference compared to the other topologies for all lengths. It has even about the same amount of interference as its equivalent Jellyfish topology. In contrast to all the other topologies, Cascade Dragonfly has two groups of node quadruples: the ones with little interference and the other group with high interference. Due to the second group of nodes with high interference, Cascade Dragonfly has the most interference of all the topologies. The most extreme difference when comparing to its equivalent Jellyfish, can be seen for $l$ = 4 and $l$ = 5. While the Jellyfish graph has at most interference 3 for any node quadruple, there are groups of nodes with interference values $>$ 15.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 52. Interference plots for Megafly with different gap sizes and Dragonfly of length $\leq$ 5. In general, Megafly has less interference than Dragonfly for length $l\leq$ 3. Only MF8 has a few outlying node pairs with a high interference at $l$ = 3. All Megafly variants have most node pairs with interference between 0 and 3. For $l\geq$ 4, all variants start to have a few $\approx$ 10 outliers with higher interference. This cannot be seen in the Dragonfly topology.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 53. Two-dimensional detailed interference plots for Megafly with different gap sizes and Dragonfly. For lengths of 2 and 3, only a few components contribute to the total interference in the Megafly networks. This changes for $l>$ 3, where almost all values for $c_{l}(\{a,c\},\{b\})+c_{l}(\{a,c\},\{d\})$ are possible. The same cannot be said for the analysis of the Dragonfly graph. For all lengths only a subset of possible values for the components contribute the total interference. This is also true for all Jellyfish equivalent graphs.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 54. Interference plots for k-ary n-trees for different values for k, Fat Trees and Fat Trees2x. For all base topologies, most node pairs have an interference of zero. Both Fat Tree variants have a few outliers that have a higher interference for length 2 and 3. The k-ary 3-tree has these outliers also, although only for $l\geq$ 4. There are also more outlying node pairs than in the Fat Tree variants.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 55. Two-dimensional detailed interference plots for the k-ary 3-tree, Fat Tree and Fat Tree2x. For both components, there exist only specific values they can get. Since values in the diagonal cancel each other out, we can see that there is only one peak next to interference zero for all base topologies.
[…]
[Main findings are in the caption.]Main findings are in the caption.
Figure 56. Interference plots for xGFTs of different variants and k-ary n-tree. Similar to before, all topologies have interference of zero for most node pairs. We can also see the peaks at a specific interference value. xGFT8S has the lowest interference of all of them. It has only interference higher than zero for $l>$ 3 and the peaks have only interference of 3. xGFT8 has only an interference $>$ 0 at length 2 and 3.
[…]
[Main findings are in the caption.]Main findings are in the caption.

### v4

@@
 EvalNet: A Practical Toolchain for Generation and Analysis of Extreme-Scale Interconnects
-Abstract.
+Abstract
 The diversity of communication paths in a network—especially non-minimal paths—is a key enabler of performance at extreme scales.
-We present EvalNet, a toolchain for scalable generation and analysis over 25 important network topologies, such as Slim Fly, PolarFly, and Orthogonal Fat Trees, with a strong focus on path diversity metrics.
+We present EvalNet, a toolchain for scalable generation and analysis of over 25 important network topologies, such as Slim Fly, PolarFly, and Orthogonal Fat Trees, with a strong focus on path diversity metrics.
 EvalNet provides an extensive and fine-grained analysis of shortest and non-shortest paths, including their multiplicities, lengths, and interference.
@@
 Using EvalNet, we deliver the widest and most comprehensive path diversity study to date, demonstrating how path diversity underpins throughput and scalability, and facilitating progress towards new frontiers in extreme-scale network design.
-copyright: noneccs: Networks Network performance evaluationccs: Networks Network performance analysisccs: Networks Network experimentationccs: Networks Data center networksccs: Networks Packet-switching networks
-1. INTRODUCTION
-The interconnection network plays an important role in today’s large-scale computing systems (Besta and Hoefler, 2014; Niranjan Mysore et al., 2009; Handley et al., 2017; Valadarsky et al., 2015; Besta et al., 2021; Di Girolamo et al., 2019; Copik et al., 2021; Hoefler et al., 2020; Di Girolamo et al., 2022).
-Large networks with tens of thousands of nodes are deployed in warehouse-sized HPC and data centers (Fu et al., 2016; Liao et al., 2014).
+I INTRODUCTION
+The interconnection network plays an important role in today’s large-scale computing systems [15, 78, 42, 100, 13, 32, 28, 45, 31].
+Large networks with tens of thousands of nodes are deployed in warehouse-sized HPC and data centers [37, 72].
 Future supercomputers as well as mega data centers will require even larger scales with hundreds of thousands of servers.
@@
 However, the evaluation of all these aspects is challenging.
-First, there are many networks that must be considered for comparison when designing a new topology, for example Fat Tree (Leiserson et al., 1996), Hypercube (Dally and Towles, 2003), Torus (Dally and Towles, 2003), Dragonfly (Kim et al., 2008), Slim Fly (Besta and Hoefler, 2014), Xpander (Ahn et al., 2009), or Jellyfish (Singla et al., 2012) – to name only a few.
+First, there are many networks that must be considered for comparison when designing a new topology, for example Fat Tree [69], Hypercube [29], Torus [29], Dragonfly [63], Slim Fly [15], Xpander [1], or Jellyfish [94] – to name only a few.
 On top of that, many of these baseline networks, for example Xpander or Orthogonal Fat Trees, do not have straightforward construction algorithms, making it challenging and time-consuming to compare to these baselines.
 Similarly, there are many performance measures that should be considered for overall evaluation, and these measures often require different tools.
-A core challenge in evaluating interconnects is the analysis of path diversity, which has emerged as a critical enabler of performance and scalability in modern networks (Besta et al., 2020).
+A core challenge in evaluating interconnects is the analysis of path diversity, which has emerged as a critical enabler of performance and scalability in modern networks [17].
 As networks grow to extreme sizes, there is a growing need for fine-grained and scalable tools that can assess not only the existence of diverse paths but also their structure, interference, and utility for multipath routing.
@@
 Throughput is ideally evaluated using packet-level simulators, which offer high accuracy but often do not scale to large topologies.
-Flow-based simulations scale more easily, but they tend to produce less accurate predictions (Casanova et al., 2008).
+Flow-based simulations scale more easily, but they tend to produce less accurate predictions [23].
 Furthermore, additional aspects like construction cost and power consumption must be analyzed.
 Unfortunately, the tools supporting these various metrics often rely on different formats for specifying networks, which significantly complicates integrated and efficient evaluation.
-To address all these issues, we propose EvalNet: a publicly available toolchain that facilitates the integrated and comprehensive evaluation of networks.
+To address all these issues, we propose EvalNet: a publicly available toolchain that facilitates the integrated and comprehensive evaluation of networks (contribution 1).
 EvalNet consists of several modules, some of which were developed from scratch, while others are interfaces to existing tools which facilitate seamless operation.
 First, EvalNet enables generation of a very large number of network topologies.
-We cover the state-of-the-art class of low-diameter topologies developed by optimizing the network structure towards the Moore Bound (Slim Fly (Besta and Hoefler, 2014), PolarFly (Lakhotia et al., 2022), BundleFly (Lei et al., 2020), Xpander (Valadarsky et al., 2015)), Dragonfly topologies (fully-connected Dragonfly (Kim et al., 2008), Cascade Dragonfly (Faanes et al., 2012)), random networks (Jellyfish (Singla et al., 2012)), torus related topologies (Mesh, Torus, Hypercube (Dally and Towles, 2003), Express Mesh (Jain et al., 2017), HyperX (Ahn et al., 2009), Flattened Butterfly (Kim et al., 2007), Tofu (Ajima et al., 2018, 2012)), indirect topologies (Fat Tree (Leiserson et al., 1996), Multi-Layer Full-Mesh (Kathareios et al., 2015), Orthogonal Fat Tree (Kathareios et al., 2015), eXtended Generalized Fat Tree (Ohring et al., 1995), k-ary n-tree (Petrini and Vanneschi, 1997)), and Kautz networks (Undirected Kautz (Li et al., 2004), Arrangement Network (Day and Tripathi, 1992)).
-Figure 1. An overview of EvalNet.
-[…]
-[Figure is fully described in the text.]Figure is fully described in the text.
-Second, we observe that the high performance of network workloads is becoming increasingly dependent on effectively harnessing various forms of path diversity, particularly through non-minimal multipathing (Besta et al., 2020).
+We cover the state-of-the-art class of low-diameter topologies developed by optimizing the network structure towards the Moore Bound (Slim Fly [15], PolarFly [65], BundleFly [68], Xpander [100]), Dragonfly topologies (fully-connected [63], Cascade [34]), random networks (Jellyfish [94]), torus-related topologies (Mesh, Torus, Hypercube [29], Express Mesh [54], HyperX [1], Flattened Butterfly [64], Tofu [3, 4]), indirect topologies (Fat Tree [69], Multi-Layer Full-Mesh [60], Orthogonal Fat Tree [60], eXtended Generalized Fat Tree [79], k-ary n-tree [84]), and Kautz networks (Undirected Kautz [70], Arrangement Network [30]).
+Figure 1: An overview of EvalNet.
+Second, we observe that the high performance of network workloads is becoming increasingly dependent on effectively harnessing various forms of path diversity, particularly through non-minimal multipathing [17, 65].
 This is especially critical in state-of-the-art low-diameter topologies, such as Slim Fly and PolarFly, which typically offer limited shortest-path diversity but compensate with a rich structure of non-minimal paths.
 In such networks, the ability to balance load and avoid congestion hinges on the capacity to exploit this non-shortest-path structure.
-To address this need, EvalNet provides a comprehensive suite of fine-grained tools for analyzing path diversity.
+To address this need, EvalNet provides a comprehensive suite of fine-grained tools for analyzing path diversity (contribution 2).
 It supports the computation of shortest and non-shortest path multiplicities, detailed distributions of path lengths, and interference-aware metrics that capture how routes interact with each other.
-EvalNet also includes visualization and measurement tools that expose the exact bandwidth and multipathing potential between every pair of routers, enabling both coarse-grained insights and fine-grained optimization of routing and scheduling schemes (Lakhotia et al., 2023).
-All path diversity analyses are designed to scale to very large topologies, supporting networks with more than 100,000 servers, thus making them practical for evaluating future-generation systems.
-Moreover, EvalNet seamlessly interfaces with different tools for network simulation.
-This includes the flow-level evaluation tool TopoBench (Jyothi et al., 2016) that enables approximating throughput for arbitrary traffic patterns and comes with a specially designed worst-case scenario that puts particularly large stress on the interconnect.
-EvalNet also harnesses two complementary packet-level simulators (htsim (Raiciu et al., 2010) and Booksim2 (Jiang et al., 2013)), which target, respectively, lossy and lossless architectures [^1] EvalNet does not compete with existing simulators as it does not introduce a new simulator.
-Instead, it tunes two of these simulators and harnesses them within its full comprehensive pipeline..
-We tune these simulators to support large-scale evaluations on low-cost hardware, achieving simulations of more than 1 million endpoint networks on a commodity PC.
-Finally, EvalNet also comes with a construction cost and power consumption assessment module.
+EvalNet also includes visualization and measurement tools that expose the exact bandwidth and multipathing potential between every pair of routers, enabling both coarse-grained insights and fine-grained optimization of routing and scheduling schemes [66].
+All path diversity analyses are designed to scale to very large topologies, supporting networks with more than 1 million servers, thus making them practical for evaluating future systems.
+We design EvalNet to enable simulations of massive networks with different architectures even on low-cost commodity hardware (contribution 3).
+For this, EvalNet harnesses two complementary packet-level simulators (htsim [91] and Booksim2 [56]), which target, respectively, lossy and lossless architectures (EvalNet does not compete with existing simulators as it does not introduce a new simulator.
+Instead, it tunes two of these simulators and harnesses them within its full comprehensive pipeline).
+We tune these simulators to support large-scale evaluations on low-cost hardware, achieving simulations of networks with more than 1 million endpoints on a commodity PC with just 16GB of RAM without impacting the simulation accuracy.
+On top of that, EvalNet also seamlessly interfaces with the flow-level evaluation tool TopoBench [57] that enables approximating throughput for arbitrary traffic patterns and comes with a specially designed worst-case scenario that puts particularly large stress on the interconnect.
+Finally, EvalNet also provides a construction cost and power consumption assessment module.
 The user can parametrize this module with arbitrary cost and power models for both routers and cables, considering arbitrary vendors and architectures such as Ethernet or InfiniBand.
-Finally, we use EvalNet to conduct the widest-ever comparative analysis of interconnect topologies, focusing on the path diversity measures, but also combining them with classical metrics that include latency, global bandwidth, construction cost, and power consumption.
+Finally, we use EvalNet to conduct the widest-ever comparative analysis of interconnect topologies (contribution 4), focusing on the path diversity measures, but also combining them with classical metrics that include latency, global bandwidth, construction cost, and power consumption.
 We investigate path diversity in-depth because other measures (latency, bandwidth under flow/packet simulations, fault-tolerance) have already been extensively researched.
 However, path diversity has only been analyzed in a limited scope.
-We provide a total of 55 rich plots for path diversity metrics, combined with observations, analyses, and insights that could be used to develop better routing strategies.
-2. EVALNET TOOLCHAIN
-We illustrate the overall structure of EvalNet in Figure 1.
-First, the user selects a topology to be generated; it can be any of the networks described in Section 3; the user can also provide a script generating their own interconnect.
-The user can then use a module for analyzing the path diversity of a network, which can subsequently be used to enhance routing schemes in simulations (an overview is in Section 4.1, details and evaluation are in Section 5).
-The generated networks can also be evaluated using either flow or packet-based simulations (overview is in Section 4.2, details and evaluation are in Section 6 and 7, respectively).
-These simulations are conducted by external tools (TopoBench (Jyothi et al., 2016) for flow-based and htsim (Raiciu et al., 2010) as well as Booksim2 (Jiang et al., 2013) for packet-based simulations of lossy and lossless networks, respectively); we provide input/output conversion routines to integrate these tools seamlessly.
-Finally, the user can use our cost and power models to assess the construction cost or static power consumption (Section 4.3).
-We now proceed to outline these factors (Sections 34) and then detail each of them and provide extensive evaluations (Sections 57).
-We will finalize with a comprehensive summary comparing all considered interconnects and measures (Section 8).
-3. RELEVANT NETWORK TOPOLOGIES
-We first describe network topologies included in EvalNet – see also Figure 2. The most important symbols are listed in Table 1.
+We provide a total of 55 plots rich in path diversity information based on 4 different metrics, combined with observations, analyses, and insights that can be used to develop better routing strategies.
+Figure 2: An overview of network topologies offered by default in EvalNet. New topologies can be straightforwardly added to the pipeline.
+II EVALNET TOOLCHAIN
+Figure 1 illustrates the overall structure of EvalNet.
+First, the user selects a topology to be generated; it can be any of the networks described in Section III; the user can also provide a script generating their own interconnect.
+The user can then use a module for analyzing the path diversity of a network, which can subsequently be used to enhance routing schemes in simulations (an overview is in Section IV-A, details and evaluation are in Section V).
+The generated networks can also be evaluated using either flow or packet-based simulations (overview is in Section IV-B, details and evaluation are in Section VI and VII, respectively).
+These simulations are conducted by external tools (TopoBench [57] for flow-based and htsim [91] as well as Booksim2 [56] for packet-based simulations of lossy and lossless networks, respectively); we provide input/output conversion routines to integrate these tools seamlessly.
+Finally, the user can use our cost and power models to assess the construction cost or static power consumption (Section IV-C).
+We now proceed to outline these factors (Sections IIIIV) and then detail each of them and provide extensive evaluations (Sections VVII).
+We will conclude with a comprehensive summary comparing all considered interconnects and measures (Section VIII).
+III RELEVANT NETWORK TOPOLOGIES
+We first describe network topologies included in EvalNet – see also Figure 2. The most important symbols are listed in Table I.
+TABLE I: The most important symbols used in this work.
 $V,E$ Sets of vertices/edges (routers/links, $V=\{0,\dots,N_{r}-1\}$).
-$N,N_{r}$endpoints and #routers in the network ($N_{r}=|V|$).
-$p,k^{\prime}$endpoints attached to a router, #channels to other routers.
+$N,N_{r}$ #endpoints and #routers in the network ($N_{r}=|V|$).
+$p,k^{\prime}$ #endpoints attached to a router, #channels to other routers.
 $k$ Router radix ($k=k^{\prime}+p$).
@@
 $\lambda,v$ Flow arrival rate [flows/s] and the flow volume [bytes].
-Table 1. The most important symbols used in this work.
-Figure 2. An overview of network topologies offered by default in EvalNet. New topologies can be straightforwardly added to the pipeline.
-[…]
-[Figure is fully described in the text.]Figure is fully described in the text.
-3.1. Network Model
+III-A Network Model
 We model an interconnection network as an undirected graph $G=(V,E)$; $V$ and $E$ are sets of vertices and edges.
@@
 $\lambda$ and $v$ are the flow arrival rate [flows/s] and the flow volume [bytes] of a used workload.
-While EvalNet supports arbitrary configurations of $p$, $k$, and $k^{\prime}$, in the experiments, we usually set $p=\frac{k^{\prime}}{d}$, as it maximizes throughput while minimizing congestion and network cost (assuming a random uniform traffic pattern) (Besta et al., 2020).
+While EvalNet supports arbitrary configurations of $p$, $k$, and $k^{\prime}$, in the experiments, we usually set $p=\frac{k^{\prime}}{d}$, as it maximizes throughput while minimizing congestion and network cost (assuming a random uniform traffic pattern) [17].
 We also usually select network radix $k^{\prime}$ and router count $N_{r}$ so that, for a fixed $N$, the compared topologies use similar amounts of networking hardware and thus have similar construction costs.
 We assume router-based topologies hereafter, i.e., no direct channels between any two endpoints.
-3.2. Moore Bound Topologies
-A recent class of low-diameter topologies explicitly optimizes the network structure towards the Moore Bound (Besta and Hoefler, 2014).
+Figure 3: An overview of evaluation measures of network topologies offered by default in EvalNet. New measures can be straightforwardly added to the pipeline.
+III-B Moore Bound Topologies
+A recent class of low-diameter topologies explicitly optimizes the network structure towards the Moore Bound [15].
 This maximizes the number of routers and endpoints for a fixed diameter and radix, minimizing cost and power consumption.
 These topologies are important as they have achieved new frontiers in performance, cost, and power consumption.
-Slim Fly (SF) (Besta and Hoefler, 2014), PolarFly (PF) (Lakhotia et al., 2022), and PolarStar (Lakhotia et al., 2024) belong to this class, they are based on the underlying McKay-Miller-Širáň and Erdős-Rényi polar graphs, respectively.
-Spectralfly (SpF) (Young et al., 2022) is based on the Ramanujan graph construction of Lubotzky, Phillips, and Sarnak (LPS).
-Moreover, a $\tau$-regular Xpander (XP) (Valadarsky et al., 2015) constructions starts with a complete graph of $\tau+1$ vertices, and is then repeatedly “lifted” ($\frac{N_{r}}{\tau+1}$ times) to obtain the final graph with good expansion properties.
-Finally, BundleFly (BF) (Lei et al., 2020) is using SF together with a Paley graph.
+Slim Fly (SF) [15], PolarFly (PF) [65], and PolarStar [67] belong to this class, they are based on the underlying McKay-Miller-Širáň and Erdős-Rényi polar graphs, respectively.
+Spectralfly (SpF) [106] is based on the Ramanujan graph construction of Lubotzky, Phillips, and Sarnak (LPS).
+Moreover, a $\tau$-regular Xpander (XP) [100] construction starts with a complete graph of $\tau+1$ vertices, and is then repeatedly “lifted” ($\frac{N_{r}}{\tau+1}$ times) to obtain the final graph with good expansion properties.
+Finally, BundleFly (BF) [68] is using SF together with a Paley graph.
 The routers in both graphs are linked via a multi-star product based on a bijection.
-3.3. Dragonfly Topologies
-The Fully-Connected Dragonfly (DF) (Kim et al., 2008) is comprised of $g$ groups of nodes internally interconnected in an all-to-all fashion.
+III-C Dragonfly Topologies
+The Fully-Connected Dragonfly (DF) [63] is comprised of $g$ groups of nodes internally interconnected in an all-to-all fashion.
 The topology among groups is an all-to-all, too.
-A Cascade Dragonfly (C-DF) (Faanes et al., 2012) alters the original Dragonfly by assuming a 16x6 2D HyperX for intra-group and an all-to-all graph for inter-group connectivity.
-Megafly (MF) (Flajslik et al., 2018) is another Dragonfly variant.
+A Cascade Dragonfly (C-DF) [34] alters the original Dragonfly by assuming a 16x6 2D HyperX for intra-group and an all-to-all graph for inter-group connectivity.
+Megafly (MF) [36] is another Dragonfly variant.
 These networks are relevant as they come with good performance and cost, albeit achieving worse values for these metrics than low-diameter networks, but have been found to be easier to cable.
-They also offer relatively easy scheduling: their modular structure facilitates achieving data locality for workloads (Tate et al., 2014).
-3.4. Random Topologies
+They also offer relatively easy scheduling: their modular structure facilitates achieving data locality for workloads [98].
+III-D Random Topologies
 Fully random networks are infeasible due to the unbounded radix.
-The Jellyfish (Singla et al., 2012) is a random graph with a pre-selected router radix $k$.
-We support these networks as they come with advantageous performance, cost, and power consumption (Singla et al., 2012; Besta et al., 2020).
-Figure 3. An overview of evaluation measures of network topologies offered by default in EvalNet. New measures can be straightforwardly added to the pipeline.
-[…]
-[Overview of evaluation measures for network topologies.]The figure provides an overview of the structure of the Evaluating Network Topologies chapter as well as the considered metrics.
-The chapter consists of path diversity measures (4.1), cost and power (4.3) as well as throughput and latency (4.2) subsections.
-The last subsection considers throughput (saturation for a given workload), maximum achievable throughput (MAT), network latency and average path length.
-The second subsection looks at the construction cost and power consumption.
-The path diversity subsection investigates counts for minimal and non-minimal paths as well as path interference and total network load (TNL).
-For the minimal and non-minial paths specifically lengths and multiplicity are considered.
-3.5. Torus Related Topologies
+The Jellyfish [94] is a random graph with a pre-selected router radix $k$.
+We support these networks as they come with advantageous performance, cost, and power consumption [94, 17].
+III-E Torus-Related Topologies
 We consider traditional Mesh (MH), Torus (TR), and Hypercube (HC) topologies.
-We also incorporate their more recent enhancements: HyperX (HX) (Ahn et al., 2009) (each router in $n$-D grid connects to all other routers within the same dimension), Express Meshes (EM) (Jain et al., 2017) (cut-down versions of the HyperX), Flattened Butterfly (FBF) (Kim et al., 2007) (a 2D variation of HyperX), and Tofu (TF) (Ajima et al., 2018, 2012) (a 6D torus variant with inner 3D mesh/tori networks embedded as virtual routers).
+We also incorporate their more recent enhancements: HyperX (HX) [1] (each router in $n$-D grid connects to all other routers within the same dimension), Express Meshes (EM) [54] (cut-down versions of the HyperX), Flattened Butterfly (FBF) [64] (a 2D variation of HyperX), and Tofu (TF) [3, 4] (a 6D torus variant with inner 3D mesh/tori networks embedded as virtual routers).
 These networks have been widely deployed across different settings and thus are considered in EvalNet.
-3.6. Indirect Topologies
+III-F Indirect Topologies
 In indirect networks, some routers attach endpoints while other routers only forward packets.
-A Fat Tree (FT) (Leiserson et al., 1996) is a binary tree, with either higher bandwidth links towards the root or multiple parallel links towards the root to avoid communication bottlenecks.
-It is one of the most widely researched and used network topologies (Katta et al., 2016b; Vanini et al., 2017; Alizadeh et al., 2014; Katta et al., 2016a; He et al., 2015; Zhang et al., 2017; Huang et al., 2020; Alasmar et al., 2018; Montazeri et al., 2018; Hwang et al., 2014; Lu, 2016; Vamanan et al., 2012; Alizadeh et al., 2013; Bai et al., 2014; Raiciu et al., 2011; Handley et al., 2017; Zhuo et al., 2016; He et al., 2016; Alizadeh et al., 2010; Cardwell et al., 2017; Mittal et al., 2015; Li and Pan, 2013; Perry et al., 2014; Olteanu and Raiciu, 2016; Park et al., 2019; Benet et al., 2018; Lu et al., 2018; Widmer et al., 2001; Callegari et al., 2014).
-The k-ary n-tree (AT) (Petrini and Vanneschi, 1997) avoids infeasible high-radix roots of FT by using the same number of routers per tree-level and connecting them with increasing strides towards the top.
-An eXtended Generalized Fat Tree (xGFT) (Ohring et al., 1995) is a k-ary n-tree variant which allows reducing deployment cost.
-Orthogonal Fat-Trees (OFT) (Kathareios et al., 2015) are based on k-ary 2-trees; the wiring pattern between the two levels is modified to increase scalability.
-Finally, a Multi-Layer Full-Mesh (MLFM) (Kathareios et al., 2015) consists of multiple 2D meshes (incl.
+A Fat Tree (FT) [69] is a binary tree, with either higher bandwidth links towards the root or multiple parallel links towards the root to avoid communication bottlenecks.
+It is one of the most widely researched and used network topologies [62, 102, 6, 61, 43, 109, 48, 5, 76, 49, 73, 101, 8, 9, 90, 42, 111, 44, 7, 22, 75, 71, 83, 80, 81, 11, 74, 104, 21].
+The k-ary n-tree (AT) [84] avoids infeasible high-radix roots of FT by using the same number of routers per tree-level and connecting them with increasing strides towards the top.
+An eXtended Generalized Fat Tree (xGFT) [79] is a k-ary n-tree variant which allows reducing deployment cost.
+Orthogonal Fat-Trees (OFT) [60] are based on k-ary 2-trees; the wiring pattern between the two levels is modified to increase scalability.
+Finally, a Multi-Layer Full-Mesh (MLFM) [60] consists of multiple 2D meshes (incl.
 diagonal connections) stacked together, but instead of direct links it uses global routers to link meshes.
 EvalNet also supports indirect networks as they offer very high bandwidth; FTs are also widely deployed.
-3.7. Kautz Topologies
-Both Undirected Kautz (KA) (Li et al., 2004) and Arrangement Network (Day and Tripathi, 1992) have strings assigned to each router.
+III-G Kautz Topologies
+Both Undirected Kautz (KA) [70] and Arrangement Network [30] have strings assigned to each router.
 Then, links between routers are determined by similarities of the corresponding router strings.
-Kautz networks offer high resilience properties and offer an interesting alternative to other classes of interconnects.
-3.8. Other Topologies
-EvalNet can be easily extended to model and evaluate interconnects from other domains, such as on-chip networks between different cores (Ahn et al., 2015; Besta et al., 2018), chiplets (Iff et al., 2023a, 2025, b), or FPGA interconnects (Cai et al., 2015), due to its modular and flexible architecture.
-The toolchain’s seamless integration with external simulators for packet-level and flow-level analysis ensures that new topologies, such as those designed for heterogeneous (Chang et al., 2012) or spatial computing environments (Gianinazzi et al., 2023; Baumann et al., 2024), can be evaluated with minimal effort, making EvalNet a versatile tool for broader interconnect research.
-4. EVALUATING NETWORK TOPOLOGIES
+Kautz networks offer high resilience properties and are an interesting alternative to other classes of interconnects.
+III-H Other Topologies
+EvalNet can be easily extended to model and evaluate interconnects from other domains, such as on-chip networks between different cores [2, 14], chiplets [50, 52, 51], or FPGA interconnects [20], due to its modular and flexible architecture.
+The toolchain’s seamless integration with external simulators for packet-level and flow-level analysis ensures that new topologies, such as those designed for heterogeneous [24] or spatial computing environments [40, 10], can be evaluated with minimal effort, making EvalNet a versatile tool for broader interconnect research.
+IV EVALUATING NETWORK TOPOLOGIES
 We next describe the considered measures used to evaluate a network topology.
-We analyzed a large body of related work to provide a comprehensive overview (Besta et al., 2021; Kim et al., 2008, 2007; Besta and Hoefler, 2014; Prisacari et al., 2013b; Yuan et al., 2014, 2013; Prisacari et al., 2014a, b; Kathareios et al., 2015; Prisacari et al., 2015; Chen et al., 2016; Prisacari et al., 2013a; Karacali et al., 2018; Sehery and Clancy, 2017; Kassing et al., 2017; Dally and Towles, 2003).
+We analyzed a large body of related work to provide a comprehensive overview [13, 63, 64, 15, 89, 107, 108, 85, 86, 60, 87, 25, 88, 58, 93, 59, 29].
 We distinguish five classes of measures, used to analyze five aspects of a network: path diversity, bandwidth, latency, construction cost, and power consumption.
 All these metrics are also listed in Figure 3.
-4.1. Path Diversity Measures
+IV-A Path Diversity Measures
 Recent studies in high-performance interconnect design indicate that harnessing different forms of path diversity, especially non-minimal multipathing, is a key performance enabler, especially for state-of-the-art low-diameter networks such as SF or PF.
-This is because these networks – while ensuring low latency, high throughput, as well as low construction cost and power consumption – have only one shortest path between almost all router pairs (Besta et al., 2020, 2021).
-Hence, one cannot use traditional load-balancing mechanisms such as equal-cost multi-path routing (ECMP) (Hopps, 2000) for more performance, and must resort to multipath routing over non-minimal paths (Besta et al., 2020).
+This is because these networks – while ensuring low latency, high throughput, as well as low construction cost and power consumption – have only one shortest path between almost all router pairs [17, 13].
+Hence, one cannot use traditional load-balancing mechanisms such as equal-cost multi-path routing (ECMP) [47] for more performance, and must resort to multipath routing over non-minimal paths [17].
 To address this, EvalNet supports analyzing all recently proposed measures for path diversity.
@@
 Finally, EvalNet enables computing the total network load (TNL), which is an upper bound on the number of flows that a network can maintain without congestion; TNL constitutes the maximum supply of path diversity offered by a given topology.
-We define all these measures in more detail in Section 5 while analyzing their values for different topologies.
+We define all these measures in more detail in Section V while analyzing their values for different topologies.
 Many of the above measures can be analyzed at different levels of detail.
@@
 EvalNet supports both options.
-Figure 4.
-(left) Histograms of shortest path lengths $l_{\text{min}}$.
+Figure 4: (left) Histograms of shortest path lengths $l_{\text{min}}$.
 (right) Histograms of shortest path multiplicities $n_{\text{min}}$.
 Considered networks: DF, SF, and two HX variants with $D\in\{2,3\}$.
-[…]
-[Histograms of shortest path lengths and multiplicities]The figure shows in two subfigures the histograms of the shortest path lengths and multiplicities for four topologies: Dragonfly, Slim Fly and HyperX (diameter of 2 and 3).
-Four different server sizes are illustrated: 1.000, 10.000, 100.000 and 1.000.000.
-The shortest path lengths facilitate reasoning about the expected latency of a topology with shortest-path minimum routing, for different router pairs.
-More router pairs in Slim Fly are connected with a 2-hop shortest path, than in HyperX (diameter of 2), for all topology sizes.
-This motivates more focus on making best-possible routing decisions over 2-hop paths in Slim Fly; 1-hop paths are negligible (especially for large networks).
-The shortest path multiplicities enable taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing.
-Nearly all router pairs in Slim Fly have only 1 shortest path between them, making non-minimal multipathing crucial.
-Most router pairs in HyperX have many shortest paths between them, and this ratio is growing with the number of servers, making ECMP a good option.
-4.2. Throughput & Latency Measures
+IV-B Throughput & Latency Measures
 In EvalNet, we offer seamless interfaces to established tools, facilitating assessing throughput using packet-based simulations.
@@
 Thus, latency measures the end-to-end communication delay without congestion effects in the network.
-Moreover, the interfaced simulators enable delivering the average path length, i.e., the average number of router-router links traversed between any two communicated endpoints.
+Moreover, the interfaced simulators enable delivering the average path length, i.e., the average number of inter-router links traversed between any two communicated endpoints.
 This measure is useful because it is related to the network topological structure, i.e., one can see more clearly (than with latency) the impact of a routing design decision on the number of links that must be traversed.
-4.3. Construction Cost & Power Consumption
+IV-C Construction Cost & Power Consumption
 EvalNet enables assessing the network construction cost.
@@
 This enables rudimentary comparisons between different baselines and preliminary insights into power consumption; detailed power models are notoriously hard to devise and vendors resort to their in-house infrastructures.
-5. PATH DIVERSITY ANALYSIS
+V PATH DIVERSITY ANALYSIS
 We present path diversity measures and conduct analyses in EvalNet, obtaining interesting insights about various interconnects.
-These measures are defined on pairs of routers (Besta et al., 2020).
+These measures are defined on pairs of routers [17].
 Extended results are provided in the Appendix.
 We start with basic definitions shared by all measures.
-Definition 5.1.
+Definition V.1.
 The neighborhood $h(A)$ of a set of vertices $A\subseteq V$ contains all the vertices adjacent to vertices in $A$, i.e., $h(A)=\left\{t\in V:\exists_{s\in A}(s,t)\in E\right\}.$
 Intuitively, $h(A)$ are all the routers connected to routers in set $A$.
-Definition 5.2.
+Definition V.2.
 The $l$-step neighborhood $h^{l}(A)$ (or a transitive hull) of $A\subseteq V$ contains all the vertices reachable by a path of length $l$ from some vertex in $A$.
@@
 Intuitively, $h^{l}(A)$ are all the routers reachable with $l$ hops from routers in set $A$.
-Definition 5.3.
+Definition V.3.
 For $s,t\in V$, the number of length-$l$ paths $n_{l}(s,t)$ from $s$ to $t$ is recursively defined as $n_{l}(s,t)=\sum_{u\in V}n_{l-1}(s,u)\cdot n_{1}(u,t),$ $n_{1}(s,t)=1\text{ iff }(s,t)\in E,\text{ else }0.$
-Figure 5. The analysis of non-shortest path diversity and path interference.
-[…]
-[Histograms of multiplicities of non-shortest paths and path interferences for shortest & non-shortest paths as well as the path diversity between all router pairs and 2D histograms of path interferences for shortest & non-shortest paths]The figure analyses non-shortest path diversity and path interference in four subfigures.
-The compared topologies are Dragonfly, SlimFly and HyperX (diameter of 2 and 3).
-The multiplicities of non-shortest paths facilitate reasoning about the expected latency of a topology with non-shortest path routing, for different router pairs, and offers hints on routing design.
-HyperX with a diameter of 2 has very high path diversity that is also ’stable’ (20 for all router pairs) for all higher path lengths (l=3,4,5).
-This facilitates the design of complex adaptive routing that routes around hotspots, uses complex paths, et cetera.
-The analysis of path interferences for shortest & non-shortest paths facilitates reasoning about the performance of both minimal and non-minimal routing, and of scheduling of potential processes to different servers.
-Dragonfly, HyperX with a diameter of 3, and Slim Fly have almost no interference at two hops.
-Thus, one does not have to pay much attention to how to schedule pairs of communicating jobs, assuming they are executed on servers attached to routers separated by the distance l=2 (or l=1).
-The analysis of the path diversity between all router pairs facilitates reasoning about the expected latency of a topology with non-shortest path routing, for different router pairs, and offers hints on how to route.
-The majority of router pairs are connected by a number of different edge-disjoint paths that is larger than 1/3 of all their available network ports (indicated with circles and triangles).
-Such pairs do not require a lot of attention when selecting a path to route.
-The analysis of path interferences for shortest & non-shortest paths acilitates reasoning about the performance of both minimal and non minimal routing, and of scheduling of potential processes to different servers is presented in the last subfigure as 2D histograms.
-5.1. Diversity of Shortest Paths
+Figure 5: The analysis of non-shortest path diversity and path interference.
+V-A Diversity of Shortest Paths
 First, we consider the lengths and multiplicities of shortest paths.
 Understanding the former facilitates reasoning about the expected latency of a topology with shortest-path minimum routing, for different router pairs.
-Knowing the latter, on the other hand, enables taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing (Besta et al., 2020, 2021).
-Definition 5.4.
+Knowing the latter, on the other hand, enables taking informed decisions on whether to use multipathing over shortest paths (e.g., with ECMP), or whether one has to use non-minimal multipath routing [17, 13].
+Definition V.4.
 For $s,t\in V$, the length of the shortest path $l_{\text{min}}(s,t)$ between them is
 $l_{\text{min}}(s,t)=\min\left\{i\in\mathbb{N}:t\in h^{i}(\{s\})\right\}.$
-Definition 5.5.
-The shortest path multiplicity between $s,t\in V$ counts the number of shortest paths between $s$ and $t$ and is defined as
+Definition V.5.
+The shortest path multiplicity between $s,t\in V$ is the number of shortest paths between $s$ and $t$ and is defined as
 $n_{\text{min}}(s,t)=n_{l}(s,t)\text{ with }l=l_{\text{min}}(s,t).$
@@
 For example, most pairs are connected with a shortest path having two hops.
-Specifically, more router pairs in SF are connected with a 2-hop shortest path, than in HX2, for all topology sizes.
-This motivates more focus on making best-possible routing decisions over 2-hop paths in SF; 1-hop paths are negligible (especially for large networks).
-The right plot shows that, for most router pairs, the multiplicity of DF, HX2, and SF is low ($n_{\text{min}}\leq 2$) compared to HX3.
+Specifically, more router pairs in SF are connected with a 2-hop shortest path than in HX2 for all topology sizes.
+This motivates a focus on making best-possible routing decisions over 2-hop paths in SF; 1-hop paths are negligible (especially for large networks).
+The right panel shows that, for most router pairs, the multiplicity of DF, HX2, and SF is low ($n_{\text{min}}\leq 2$) compared to HX3.
 We also observe that nearly all router pairs in SF have only 1 shortest path between them, making non-minimal multipathing crucial.
-Still, note that DF and SF host some outlier router pairs connected with large counts of shortest paths.
-Finally, most router pairs in HX have many shortest paths between them, and this ratio is growing with $N$, making ECMP a good option.
-5.2. Diversity of Non-Shortest Paths
+Still, note that DF and SF contain some outlier router pairs connected with large counts of shortest paths.
+Finally, most router pairs in HX have many shortest paths between them, and this ratio is growing with $N$, making ECMP effective.
+V-B Diversity of Non-Shortest Paths
 Next, we are interested in the number of edge disjoint paths between two vertices $s,t\in V$, that are not necessarily shortest.
@@
 For convenience, we define this measure for two groups of vertices $A,B$:
-Definition 5.6.
+Definition V.6.
 The count of edge disjoint paths $c_{l}(A,B)$ at length $l$ between sets of routers $A\subset V$, $B\subset V$ is the smallest number of edges that can be removed such that $A$ and $B$ are not connected by any path of length at most $l$.
 To obtain this count for two specific routers, we set $A=\{s\},B=\{t\}$. The above definition lets us easily derive the number of shortest edge disjoint paths between two vertices:
-Definition 5.7.
+Definition V.7.
 The count of shortest edge disjoint paths $c_{\text{min}}(s,t)$ between $s,t\in V$ is defined as
@@
 HX2 has very high path diversity that is also “stable” (20 for all router pairs) for all higher path lengths ($l=3,4,5$).
-This facilitates the design for complex adaptive routing that routes around hotspots, uses complex paths, et cetera.
+This facilitates the design for complex adaptive routing that routes around hotspots and uses complex paths.
 Next, DF has relatively large diversity of paths for $l=5$, suggesting similar insights as for HX2, but the count is 11, so less than 20 by nearly 50%.
@@
 These specific pairs are not good candidates for scheduling communication-intense jobs.
-5.3. Path Interference
+V-C Path Interference
 So far, we considered measures that only take into account paths between two vertices.
@@
 Here, paths between two router pairs $a,b$ and $c,d$ interfere if their total count $c_{l}(\{a,c\},\{b,d\})$ of edge disjoint paths at length $l$ is lower than the sum of the single destination count of edge disjoint paths $c_{l}(\{a,c\},\{b\})+c_{l}(\{a,c\},\{d\})$.
-Definition 5.8.
+Definition V.8.
 Path interference $I^{l}_{ac,bd}$ between two router pairs $a,b$ and $c,d$ is defined as
@@
 Such pairs still require care when scheduling communication-intense jobs.
-5.4. Computing Path Diversity Efficiently
-In EvalNet, we primarily use algebraic algorithms to compute path diversity metrics, as they offer better scalability and parallelizability compared to purely combinatorial approaches, particularly for large, low-diameter networks (Besta et al., 2017; Solomonik et al., 2017; Solomonik and Demmel, 2011; Solomonik et al., 2014).
-For example, for counting the number of paths of fixed length between all node pairs, we use adjacency matrix exponentiation, where the $\ell$-th power of the adjacency matrix $A$ gives the number of $-\ell$ paths via $(A^{\ell}){i,j}$.
+V-D Computing Path Diversity Efficiently
+In EvalNet, we primarily use algebraic algorithms to compute path diversity metrics, as they offer better scalability and parallelizability compared to purely combinatorial approaches, particularly for large, low-diameter networks [16, 95, 97, 96].
+For example, for counting the number of paths of fixed length between all node pairs, we use adjacency matrix exponentiation, where the $\ell$-th power of the adjacency matrix $A$ gives the number of length-$\ell$ paths via $(A^{\ell}){i,j}$.
 This computation is efficient for small diameters and can be adapted for next-hop tracking using semiring-based variants.
 For length-restricted connectivity, we extend the algebraic path propagation algorithm of Cheung et al.
-(Cheung et al., 2011).
+[26].
 For other metrics, we also explore max-flow-based approaches such as Push-Relabel and Ford-Fulkerson with BFS path constraints (used for sampled length-limited disjoint path analysis), as well as classic Gomory-Hu trees for all-pairs edge connectivity.
 While these combinatorial algorithms serve as correctness baselines or targeted tools for small-scale samples, EvalNet’s core design centers on the algebraic methods due to their efficiency, extensibility, and compatibility with parallel computation.
-Figure 6.
-(left) Heatmap showing the Maximum Achievable Throughout (MAT), obtained with TopoBench, for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern.
+Figure 6: (left) Heatmap showing the Maximum Achievable Throughput (MAT), obtained with TopoBench, for a SF with 10k endpoints and a variant of FatPaths routing with a worst-case traffic pattern.
 (right) Packet-based simulation runtimes [cycles] with Booksim2, for networks of 50k endpoints, for two injection rates, 0.01 and 0.5.
@@
 For load=0.5, the network global bandwidth starts to play an important role; here, DF has higher latency than FT3 because it has much lower global bandwidth.
-[…]
-[Figure is fully described in the text]Figure is fully described in the text.
-6. FLOW-LEVEL SIMULATIONS
+VI FLOW-LEVEL SIMULATIONS
 Flow-level simulators enable evaluation of large-scale systems, at a cost of a coarse design that may hinder realistic insights into performance.
@@
 Without any additional constraints, this leads to constrained max-flow solutions in the topology graph.
-Other constrains (e.g., on the routes that can be taken by specific flows) lead to more realistic upper-bounds for performance of various routing schemes on specific topologies.
+Other constraints (e.g., on the routes that can be taken by specific flows) lead to more realistic upper-bounds for performance of various routing schemes on specific topologies.
 However, since these models do not represent time accurately, they are unable to estimate latency or provide realistic insights into the impact of routing choices on flow control.
 As such, they are a useful tool for initial insights when designing a network.
-For flow simulations, we provide an interface to TopoBench, an evaluation tool (Jyothi et al., 2016) that uses linear programming (LP) to derive MAT.
-Our TopoBench version supports state-of-the-art layered routing schemes such as FatPaths (Besta et al., 2020) or extended layered routing (Blach et al., 2024), built specifically for enabling non-minimal multi-path routing.
+For flow simulations, we provide an interface to TopoBench, an evaluation tool [57] that uses linear programming (LP) to derive MAT.
+Our TopoBench version supports state-of-the-art layered routing schemes such as FatPaths [17] or extended layered routing [18], built specifically for enabling non-minimal multi-path routing.
 Example results are in the left part of Figure 6, where we analyze a variant of FatPaths that offers non-minimal multipath routing.
@@
 “#R” (Y axis) as well as “#P” (X axis) are parameters of the FatPaths routing scheme; they can be tuned for a specific topology.
-Usually, the higher #R is, the better the throughput becomes (because more layers entail more potential for multipathing (Besta et al., 2020)).
+Usually, the higher #R is, the better the throughput becomes (because more layers entail more potential for multipathing [17]).
 Simultaneously, higher #R indicates more HW costs because one needs larger forwarding tables.
-Then, higher #P does not always enhance MAT (as confirmed by the results), because certain value ranges may cause larger path interference across routing layers, ultimately diminishing performance (Besta et al., 2020).
+Then, higher #P does not always enhance MAT (as confirmed by the results), because certain value ranges may cause larger path interference across routing layers, ultimately diminishing performance [17].
 Overall, for certain configurations, FatPaths on SF achieves very high MAT.
-7. PACKET-LEVEL SIMULATIONS
+VII PACKET-LEVEL SIMULATIONS
 Packet-level simulators offer a more detailed packet-based model of the network and routing.
@@
 The main simulation loop keeps the scheduled events in a priority queue and advances the simulation by dequeuing and executing successive events.
-7.1. Which Simulator to Use?
+VII-A Which Simulator to Use?
 A large number of simulators exist, and it is challenging to select the one best suited for one’s purposes.
 We analyze a broad selection of available simulators, and show that none scale to the desired sizes of hundreds of thousands of servers.
-The overview is in Table 2.
+The overview is in Table II.
 We analyzed their codebases, investigating which (1) facilitate modifications that could enhance scalability, (2) are easy to use and extend without requiring distributed-memory infrastructure support, and (3) cover a large number of protocols and settings.
-Based on that analysis, we selected two simulators: htsim (Raiciu et al., 2010) for lossy TCP/IP stacks and Booksim2 (Jiang et al., 2013) for lossless HPC architectures.
+Based on that analysis, we selected two simulators: htsim [91] for lossy TCP/IP stacks and Booksim2 [56] for lossless HPC architectures.
 Other simulators do not meet these criteria.
 For example, SST and ROSS natively target HPC clusters and we concluded that they are relatively complex to set up, use, and extend on low-cost PCs.
-htsim The htsim simulator (Raiciu et al., 2010) provides a lightweight infrastructure for lossy networks.
+htsim The htsim simulator [91] provides a lightweight infrastructure for lossy networks.
 It only models the transport layer: there is no model for links or routers.
@@
 EvalNet offers a seamless interface to htsim.
-We select htsim as it targets TCP stacks; it supports by default modern TCP-based schemes such as NDP (Handley et al., 2017), and it can be easily extended with new routing schemes.
+We select htsim as it targets TCP stacks; it supports by default modern TCP-based schemes such as NDP [42], and it can be easily extended with new routing schemes.
 Booksim2 Booksim2 targets lossless networks.
@@
 As with htsim, EvalNet offers modules that enable seamless evaluation of any generated topologies within the Booksim2 setting.
+TABLE II: The comparison of available packet-level network simulators.
+“Scalability” is the largest size of a topology (simulated with a specific simulator) that we were able to find in the literature.
+“Design” includes the details of the compute platform used for simulations in the “Scalability” column: “SM” (shared-memory design), “DM” (distributed-memory design).
 Simulator Scalability [#servers] Design
-LogGOPSim (Hoefler et al., 2017) 1024 (Hoefler et al., 2017) SM
-htsim (Raiciu et al., 2010) 128 (Raiciu et al., 2010) SM
-NS2 (Issariyakul and Hossain, 2009) $<$1,000 SM
-booksim (Jiang et al., 2013) 10,000 (Besta and Hoefler, 2014) SM
-FOGSim (Garcıa et al., 2014) 16,512 (Benito et al., 2015) SM
-OMNet++(Varga and Hornig, 2008) 100,000 (Gamer and Mayer, 2009) SM
-NS3 (of Washington NS-3 Consortium, 2024) 5,000 (Pelkey and Riley, 2011) DM
-NS4 (Fan et al., 2017) N.A. DM
-BigSim (Zheng et al., 2004) 46,656 (Jain et al., 2016) DM
-SST (Rodrigues et al., 2016) 110,592 (Groves et al., 2016) DM
-xSim (Böhm and Engelmann, 2011) 2,097,152 (Engelmann, 2014) DM
-ROSS/CODES (Cope et al., 2011) 1,000,000 (Wolfe et al., 2016) (Slim Fly), 50,000,000 (Mubarak et al., 2012) (Dragonfly) DM
+LogGOPSim [46] 1024 [46] SM
+htsim [91] 128 [91] SM
+NS2 [53] <<1,000 SM
+booksim [56] 10,000 [15] SM
+FOGSim [39] 16,512 [12] SM
+OMNet++[103] 100,000 [38] SM
+NS3 [99] 5,000 [82] DM
+NS4 [35] N.A. DM
+BigSim [110] 46,656 [55] DM
+SST [92] 110,592 [41] DM
+xSim [19] 2,097,152 [33] DM
+ROSS/CODES [27] 1,000,000 [105] (Slim Fly), 50,000,000 [77] (Dragonfly) DM
 htsim [this work] 1,000,000 SM
 Booksim2 [this work] 50,000 SM
-Table 2.
-The comparison of available packet-level network simulators.
-“Scalability” is the largest size of a topology (simulated with a specific simulator) that we were able to find in the literature.
-“Design” includes the details of the compute platform used for simulations in the “Scalability” column: “SM” (shared-memory design), “DM” (distributed-memory design).
+TABLE III: Estimation of the number of elements of a data center Slim Fly simulation.
+The numbers and thus required resources vary widely depending on a chosen configuration.
+“Range” lists parameter bounds considered in this work, in the following columns we show numbers for example configurations.
+Memory estimates are based on htsim.
+The chosen workload parameters keep the simulations feasible on a laptop (configuration details are in Section VII-C).
+1Negligible for relevant configurations.
+2Data depends on the routing scheme.
+3Objects are transient and accounted for in the flow counts.
+4Inter-router cables + server links.
+5The average over the flow size distribution, excluding re-transmissions.
 Element Range (lower–upper bounds) 10,000 servers 100,000 servers 1,000,000 servers
@@
 path (per flow) 1-10 $10^{2}$$10^{8}$ 600B 5 4,840,000 2.7GB 5 4,761,000 2.7MB 5 5,618,000 3.2GB
-packet (per flow) 10-1000 $10^{3}$$10^{10}$ (3) (5) 110 106,480,000 (3) (5) 110 104,742,000 (3) (5) 110 123,596,000 (3)
-Table 3.
-Estimation of the number of elements of a data center Slim Fly simulation.
-The numbers and thus required resources vary widely depending on a chosen configuration.
-“Range” lists parameter bounds considered in this work, in the following columns we show numbers for example configurations.
-Memory estimates are based on htsim.
-The chosen workload parameters keep the simulations feasible on a laptop (configuration details are in 7.3).
-1 Negligible for relevant configurations.
-2 Data depends on the routing scheme.
-3 Objects are transient and accounted for in the flow counts.
-4 Inter-router cables + server links.
-5 The average over the flow size distribution, excluding re-transmissions.
-7.2. Can We Simulate Large Networks on Low-Cost HW?
+packet (per flow) 10-1000 $^{3}1010$$^{10}$ (3) (5) 110 106,480,000 (3) (5) 110 104,742,000 (3) (5) 110 123,596,000 (3)
+VII-B Can We Simulate Massive Networks on Low-Cost HW?
 Next, we conduct a feasibility analysis in which we argue that million-server packet-level simulators should in theory be achievable on a low-cost PC.
@@
 Number of Simulation Elements We estimate the number of elements of a simulated topology (e.g., servers) and of a simulated workload (e.g., flows which require flow control state).
-The results of the analysis are in Table 3.
+The results of the analysis are in Table III.
 The number of network elements does not pose the most serious scalability problems.
@@
 Memory and Time Requirements The total memory usage and simulation time depend on the simulation software and on granularity, i.e., what elements of the simulated workload are explicitly stored and simulated.
-We estimate these numbers in Table 3, which also shows what may not be feasible in the largest-scale simulations.
+We estimate these numbers in Table III, which also shows what may not be feasible in the largest-scale simulations.
 Specifically, we cannot simulate large workloads on large networks, and must not store any per-packet state.
@@
 Special attention needs to be paid to monitoring and debugging code, which can easily dominate the per-flow state.
-For very large networks, the quadratically increasing size of the routing tables starts to dominate, but as Table 3 shows, this was not a problem for our simulations.
+For very large networks, the quadratically increasing size of the routing tables starts to dominate, but as Table III shows, this was not a problem for our simulations.
 The lowest amount of time is determined by the amount of packet forwarding.
@@
 Since each event likely causes at least one cache miss, we cannot expect much higher rates without distributed simulation or middle- to high-cost PC stations, which are outside the scope of this work.
-7.3. How to Simulate Really Large Networks on Low-Cost HW?
-Finally, we show how to use htsim (Raiciu et al., 2010) and Booksim2 (Jiang et al., 2013) to enable more scalability.
+VII-C How to Simulate Massive Networks on Low-Cost HW?
+Finally, we show how to use htsim and Booksim2 to enable more scalability.
 In Booksim2, the key limiting factor that we identified was processing related to single packets.
 This processing is usually related to gathering statistics, which is not relevant for the actual simulation outcomes.
-In Booksim2, eliminating this processing was the main enabler for being able to simulate networks with 10,000 endpoints and more.
+In Booksim2, eliminating this processing was the main enabler for being able to simulate networks with 10k to 1M endpoints and more.
 In htsim, on the contrary, the main bottleneck was the storage related to flows.
@@
 Another scalability obstacle was the net_paths structure present in the sample programs, a preallocated $N^{2}$ size matrix of routes, which would only be sparsely populated for permutation workloads and would dominate memory use for large networks.
-7.4. Representative Simulation Results
-We simulate networks with $\approx$10k, $\approx$100k, and $\approx$1M servers using a laptop with 16GB RAM and an Intel Core i7-8550U CPU; each run takes 1–3 hours.
-Networks are 5$\times$ oversubscribed, as in configurations studied in (Kassing et al., 2017), and include topologies larger than those used in prior large-scale SF simulations.
-Due to the cost of simulating large networks, we capture only a few milliseconds of activity—sufficient for reliable relative performance comparisons.
-We focus on flow completion time (FCT) distributions as a function of flow size, using the pFabric web search distribution (average $\approx$1MB) and injection rates of 40–60 flows/server/s.
-Results are shown as mean, 10th, and 99th percentiles.
-To mitigate artifacts from the fixed injection window, we carefully interpret short and long flows.
-Figure 7. Results of large-scale simulations.
+VII-D Representative Simulation Results
 […]
-[Comparison of the Jellyfish and Slim Fly topologies]The figure shows three different plots comparing various aspects of the Jellyfish and Slim Fly topologies.
-The first plot presents the flow completion time with similar parameters as in the previous figure: Jellyfish and Slim are identical for the 10% percentile and the mean, but Jellyfish has a smaller completion time for the 99% percentile.
-The second subfigure plots the distribution as a histogram for 10.000, 100.000 and 1.000.000 servers for long flows (200KB) and an arrival rate of 50/s.
-Jellyfish exhibits a better tail latency.
-The third plot compares the flow completion time with respect to the arrival rate (ranging from 40 to 60) of long flows and the number of servers (10.000, 100.000 and 1.000.000).
-Both topologies exhibit the same mean performance, but Jellyfish is faster for the 10% and 90% percentile.
+Figure 7: Results of large-scale simulations.
 As an example, we analyze 10k, 100k, and 1M server simulations, see Figure 7.
 The left plot shows FCT as a function of size.
-For long flows ($v>200$kB), FCT is limited by bandwidth, while for shorter flows, latency dominates.
+For long flows (v>200v>200kB), FCT is limited by bandwidth, while for shorter flows, latency dominates.
 The middle plot illustrates the distribution of FCTs for middle-size flows: here, the better tail behavior of the randomized Jellyfish topology is clearly visible.
@@
 However, they require some compromises on the quality of simulation, most importantly on the number of allowed flows and on the methodology (i.e., simulating a fixed amount of flows instead of a fixed time window after reaching a steady-state).
-8. COMPREHENSIVE ANALYSIS & TAKEAWAYS
+VIII COMPREHENSIVE ANALYSIS & TAKEAWAYS
+TABLE IV: A general comparison of different network topologies.
+Lat.: latency.
+GlB.: global bandwidth (i.e., the saturation point for the random uniform traffic pattern).
+Ext.: extensibility (i.e., how far can we extend a given concrete network with new servers?).
+Note that fixed diameter networks have lowest extensibility, because they have strict upper bounds on how many new servers can be attached; the lower the diameter is, the lower this bound is.
+Contrarily, networks such as torus can attach arbitrarily many new servers, because their diameter grows to infinity.
+C&P: construction cost and static power consumption.
+They are assessed as proportional to the total number of ports in a network, for a fixed network size NN (i.e., for fixed $N$, networks with higher radix tend to have higher cost and power consumption).
+MS: diversity of shortest paths.
+It is assessed as a weighted average of shortest path diversities at different lengths.
+MnS: diversity of non-shortest paths.
+It is assessed as a weighted average of non-shortest path diversities at different lengths.
+The battery symbols serve as a rudimentary measure of relative comparison between networks.
+[0]: worst, [1]: bad, [2]: medium, [3]: good, [4]: excellent.
+They always have a positive meaning, e.g., “ [4]” used for cost means that the cost of a given network is very low compared to other networks.
 Topology Lat. GlB. Ext. C&P MS MnS
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Slim Fly [15] [4] [4] [0] [3] [1] [4]
+PolarFly [65] [4] [4] [0] [4] [0] [4]
+Xpander [100] [3] [4] [1] [2] [1] [3]
+BundleFly [68] [3] [3] [1] [2] [1] [3]
+Dragonfly [63] [3] [3] [1] [2] [0] [2]
+Cascade Dragonfly [34] [2] [3] [1] [2] [1] [1]
+Jellyfish [94] [2] [2] [1] [2] [1] [3]
+Mesh [0] [0] [3] [0] [0] [1]
+Torus 2D [0] [0] [3] [0] [0] [1]
+Torus 3D [1] [1] [3] [1] [1] [2]
+Torus 4D [2] [2] [3] [1] [2] [3]
+Torus 5D [3] [3] [4] [2] [3] [3]
+Torus 6D [3] [3] [4] [2] [3] [4]
+Hypercube [3] [3] [4] [2] [3] [4]
+HyperX (2-dimensional) [1] [4] [4] [1] [2] [3] [4]
+HyperX (3-dimensional) [1] [3] [4] [2] [2] [3] [4]
+Express Mesh [54] [2] [3] [3] [3] [2] [3]
+Flattened Butterfly [64] [4] [4] [1] [2] [3] [4]
+Fat Tree (2-level) [69] [4] [4] [1] [3] [4] [0]
+Fat Tree (3-level) [69] [3] [2] [2] [2] [4] [0]
+k-ary n-tree [84] [3] [4] [4] [2] [4] [0]
+eXtended Generalized Fat Tree [79] [3] [3] [4] [2] [4] [0]
+Orthogonal Fat-Trees [60] [4] [4] [1] [4] [4] [0]
+Multi-Layer Full-Mesh [60] [4] [3] [1] [2] [3] [2]
+Undirected Kautz [70] [2] [2] [1] [2] [1] [2]
+Arrangement Network [30] [1] [2] [1] [2] [0] [1]
 EvalNet enables multifaceted analysis of numerous network topologies.
-We list the considered networks in Table 4, and provide an assessment of their important properties (performance, cost & power consumption, path diversity), obtained using EvalNet.
+We list the considered networks in Table IV, and provide an assessment of their important properties (performance, cost & power consumption, path diversity), obtained using EvalNet.
 The table compiles the results described in the previous sections of this paper, as well as further results illustrated in the Appendix in Figures 856.
-8.1. General Observations & Patterns
+VIII-A General Observations & Patterns
 A central observation is the critical role of path diversity (both shortest-path (MS) and non-shortest-path (MnS)) in determining a network’s performance and scalability.
@@
 EvalNet enables these tradeoffs to be quantified and visualized precisely, supporting informed design decisions based on latency, throughput, scalability, cost-efficiency, and path diversity.
-8.2. Example Detailed Insights
+VIII-B Example Detailed Insights
 We also provide a few selected more detailed insights that could be used, e.g., as a basis for enhancements of routing strategies. Full results and observations are in the Appendix.
@@
 Thus, one can tune PolarStar for either peak throughput (Paley) or tighter latency tails (BDF) based on application’s burstiness.
-9. CONCLUSION
-In this work, we introduce EvalNet, a comprehensive toolchain for the generation and in-depth analysis of extreme-scale interconnection networks, with a strong focus on path diversity as a core enabler of performance and scalability.
+IX CONCLUSION
+We introduce EvalNet, the first toolchain that enables comprehensive generation and in-depth analysis of extreme-scale interconnection networks, with a strong focus on path diversity as a core enabler of performance and scalability.
 EvalNet integrates topology generation, fine-grained path diversity analysis, performance modeling, and cost evaluation into a unified and extensible framework.
@@
 EvalNet interfaces seamlessly with established simulators and serves as a practical and versatile tool for guiding future innovations in large-scale network design in data centers and supercomputers.
-Acknowledgments
-We thank Haoran Zhao for his code contributions.
+Acknowledgment
+We thank Ankit Singla for help during the early stages and Haoran Zhao for his code contributions.
 We thank Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault and Daint machines.
 We thank Timo Schneider for help with infrastructure at SPCL.
+This paper has been assigned the LANL identification number LA-UR-26-20454.
 This project received funding from the European Research Council (Project PSAP, No.
@@
 This project was supported by JSPS KAKENHI Grant Number JP19H04119.
+We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grants no.
+PLG/2024/017103 and PLG/2025/018259.
 Appendix A Appendix
-Figure 8.
-Shortest path plot for Kautz and Arrangement graphs.
+Figure 8: Shortest path plot for Kautz and Arrangement graphs.
 Kautz graphs have lower shortest paths compared to their Jellyfish equivalent for small networks.
@@
 In comparison, Arrangement networks have a drop-off between 5 and 10 times.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 9.
-Multiplicity plot for Kautz and Arrangement graphs.
+Figure 9: Multiplicity plot for Kautz and Arrangement graphs.
 For Arrangement networks, the shortest path density first decreases, then increases.
@@
 This cannot be seen for the Jellyfish equilibrium for $K_{6,n}$.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 10.
-Shortest path plot for mesh topology of different dimensions.
+Figure 10: Shortest path plot for mesh topology of different dimensions.
 The number of shortest paths increases while $l_{min}$ increases.
@@
 Only for $l_{min}$ = 5, the Jellyfish network with higher N does not strictly have fewer amounts of shortest paths.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 11.
-Multiplicity plot for meshes of different dimensions.
+Figure 11: Multiplicity plot for meshes of different dimensions.
 As expected, the analysis looks fairly similar across all dimensions.
@@
 For bigger values of $n_{min}$, the shortest path multiplicity drops even more compared to the original mesh, when we increase the network size.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 12.
-Shortest path plot for the 4D Express Meshes with different gap sizes.
+Figure 12: Shortest path plot for the 4D Express Meshes with different gap sizes.
 We choose the lowest g, such that for each Express Mesh 4DeMESHx we have at least x additional express connections per dimension.
@@
 They increase first before they decrease for the biggest value of $l$ where there still exists a shortest path.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 13.
-Multiplicity plot for Express Meshes of dimension 4.
+Figure 13: Multiplicity plot for Express Meshes of dimension 4.
 For small and medium-sized networks, there is a gap in the multiplicity counts.
@@
 In contrast to the base Express Meshes, the Jellyfish equilibriums are well distributed along all values for $n_{min}$ for small and medium networks.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 14.
-Shortest path plot for 6D Express Meshes with different gap sizes.
+Figure 14: Shortest path plot for 6D Express Meshes with different gap sizes.
 We chose the lowest g, such that for each Express Mesh 6DeMESHx we have at least x additional express connections per dimension.
@@
 The amount of shortest paths grows as $l$ grows, before dropping down for some outliers that have a shortest path length of 5.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 15.
-Multiplicity plot for Express Meshes of dimension 6.
+Figure 15: Multiplicity plot for Express Meshes of dimension 6.
 We can see again the gap in the amount of shortest path multiplicity.
@@
 Just for $n_{min}$ = 5, the multiplicity count increases, as they have as well the most shortest paths with multiplicity $\geq$ 5.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 16.
-Shortest path plot for the Tofu topology, 6D torus and 6D mesh.
+Figure 16: Shortest path plot for the Tofu topology, 6D torus and 6D mesh.
 As we have already seen in the analysis of the meshes, the amount of shortest paths sinks as the network size gets bigger.
@@
 The same is true when we compare Tofu to the equivalent Jellyfish.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 17.
-Multiplicity plot for the Tofu network, 6D torus and 6D mesh.
+Figure 17: Multiplicity plot for the Tofu network, 6D torus and 6D mesh.
 Similar as for the meshes, all the topologies have the most multiplicity count where $n_{min}\geq$ 5.
@@
 For the large network, there is a significant drop-off and the multiplicity is about 30 times smaller for $n_{min}\geq$ 5 compared to the small or medium network.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 18.
-Shortest path plot for PolarStar compared to Dragonfly and Slim Fly.
+Figure 18: Shortest path plot for PolarStar compared to Dragonfly and Slim Fly.
 We can clearly see that all are diameter $\leq$ 3 topologies.
@@
 There is almost no difference for the analysis between PolarStar with BDF or Paley subgraphs.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 19.
-Multiplicity plot for PolarStar, Dragonfly and Slim Fly.
+Figure 19: Multiplicity plot for PolarStar, Dragonfly and Slim Fly.
 Slim Fly has the least multiple shortest paths of all topologies.
@@
 PolarStar with BDF subgraphs has only a few shortest path multiplicity count for $n_{min}$ = 3 or 4, while PolarStar with the Paley subgraphs has a well-distributed count for all values of $n_{min}$.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 20.
-Shortest path plot for Cascade Dragonfly, PolarStar and Spectralfly.
+Figure 20: Shortest path plot for Cascade Dragonfly, PolarStar and Spectralfly.
 The amount of shortest paths of length 1 and 2 decreases, if the network gets bigger, because a single node in a group can only reach 21 nodes from another group with at most 2 jumps.
@@
 Only for large network sizes, the Jellyfish graph shows no shortest paths of length 5, while the Spectralfly graph still has some.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 21.
-Multiplicity plot for the Cascade Dragonfly, PolarStar and Spectralfly networks.
+Figure 21: Multiplicity plot for the Cascade Dragonfly, PolarStar and Spectralfly networks.
 Most node pairs in the Cascade Dragonfly topology have multiple ($\geq$ 5) shortest paths.
@@
 While the equivalent Jellyfish of Spectralfly drops in multiplicity counts for higher values of $n_{min}$, for medium and large networks it increases.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 22.
-Shortest path plot for Megafly with different values for g and Dragonfly.
+Figure 22: Shortest path plot for Megafly with different values for g and Dragonfly.
 We can see, that Megafly is a diameter 3 network as well.
@@
 The Jellyfish graphs have additional shortest paths of length 1 and sometimes length 4.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 23.
-Multiplicity plot for the Dragonfly and Megafly topologies with different values for g.
+Figure 23: Multiplicity plot for the Dragonfly and Megafly topologies with different values for g.
 All topologies have mostly only one shortest path between any two nodes.
@@
 Their equivalent Jellyfish on the other hand have well-distributed shortest path multiplicities.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 24.
-Shortest path plot for Fat Trees and the k-ary 3-tree.
+Figure 24: Shortest path plot for Fat Trees and the k-ary 3-tree.
 All topologies exhibit the same patterns for the shortest paths of a given length.
 While the shortest paths of length 4 stay the same for any network size, the number of paths of length 2 decreases as the network gets bigger.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 25. Multiplicity plot for Fat Trees and the k-ary 3-tree. All networks and all sizes have a path multiplicity count of 5 or higher.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 26.
-Shortest path plot for length $l_{min}$ up to $l$ = 5 for multiple variants of xGFTs.
+Figure 25: Multiplicity plot for Fat Trees and the k-ary 3-tree. All networks and all sizes have a path multiplicity count of 5 or higher.
+Figure 26: Shortest path plot for length $l_{min}$ up to $l$ = 5 for multiple variants of xGFTs.
 There are only shortest paths of even length.
@@
 As the network size increases, the number of shortest paths at each length decreases.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 27.
-Multiplicity plot for multiple variants of xGFTs.
+Figure 27: Multiplicity plot for multiple variants of xGFTs.
 Variant 4 has the lowest amount of multiplicity count.
@@
 For large network sizes, the 8S variant has even more multiplicity than the smaller networks of that variant.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 28.
-Edge disjoint path plots for Kautz and Arrangement graphs of small sizes.
+Figure 28: Edge disjoint path plots for Kautz and Arrangement graphs of small sizes.
 The connectivity for $A_{n,4}$ is better than the connectivity for $A_{n,5}$.
@@
 Furthermore, the Kautz networks analysis looks similar to its equivalent Jellyfish.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 29.
-Edge disjoint path plots for meshes of different dimensions.
+Figure 29: Edge disjoint path plots for meshes of different dimensions.
 For all lengths $l$, the meshes have mostly zero edge disjoint paths between any two nodes.
@@
 The 4D mesh has a slightly worse path diversity than the 5D and 6D meshes.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 30.
-Edge disjoint path plots for the Tofu topology, 6D torus and 6D Express Mesh.
+Figure 30: Edge disjoint path plots for the Tofu topology, 6D torus and 6D Express Mesh.
 Similar to the diversity analysis of the meshes, all the topologies have zero path diversity for most node pairs of any given length.
@@
 Same as before, the equivalent Jellyfish graphs have a well-distributed diversity count.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 31.
-Edge disjoint path plots for Cascade Dragonfly, PolarStar and Spectralfly topologies.
+Figure 31: Edge disjoint path plots for Cascade Dragonfly, PolarStar and Spectralfly topologies.
 Both Spectralfly and Cascade Dragonfly have mostly a low diversity count.
@@
 Even for length $l$ = 5, Spectralfly has only node pairs with 4 edge disjoint paths.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 32.
-Edge disjoint path plots for Dragonfly, PolarStar and Slim Fly topologies.
+Figure 32: Edge disjoint path plots for Dragonfly, PolarStar and Slim Fly topologies.
 Slim Fly has the most path diversity by far.
@@
 Both variants of PolarStar have the same or better diversity compared to their Jellyfish equivalent for all lengths.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 33.
-Edge disjoint path plots for Dragonfly and Megafly networks with different values of g.
+Figure 33: Edge disjoint path plots for Dragonfly and Megafly networks with different values of g.
 All Megafly networks have a low path diversity of either 0 or 1 for most node pairs and all lengths.
@@
 The diversity counts of all Jellyfish equilibriums show a Gauss distribution for $l$ = 3.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 34.
-Edge disjoint path plots for the Fat Trees and k-ary 3-tree.
+Figure 34: Edge disjoint path plots for the Fat Trees and k-ary 3-tree.
 All the topologies have only specific values for the path diversity.
@@
 The k-ary 3-tree has the best path diversity among all other topologies.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 35.
-Edge disjoint path plots for the xGFTs of different variants.
+Figure 35: Edge disjoint path plots for the xGFTs of different variants.
 There exist only edge disjoint paths of specific values.
@@
 The k-ary n-tree has edge nodes with a higher diversity count than any other topology, but it has still node pairs with a diversity count of zero for $l$ = 5.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 36.
-Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 3.
+Figure 36: Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 3.
 All router pairs have a connectivity $\leq$ 75% for $l$ = 3.
@@
 This can be seen horizontally as well as vertically.
-Due to this mesh of edges, all nodes are connected and there exist only a couple of node pairs with connectivity $<$ 25%.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 37. Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 5. All router pairs have a connectivity $\geq$ 75% for $l$ = 5.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 38.
-Low connectivity plot for the PolarStar graph of length $l$ = 3.
+Due to this mesh of edges, all nodes are connected and there exist only a couple of node pairs with connectivity << 25%.
+Figure 37: Low connectivity plot for $K_{4,3}$ Kautz graph of length $l$ = 5. All router pairs have a connectivity $\geq$ 75% for $l$ = 5.
+Figure 38: Low connectivity plot for the PolarStar graph of length $l$ = 3.
 The structure graph is an ER2 graph and the subgraph is a BDF3 graph.
@@
 Every node in 0 - 7 is connected to a node in 31 - 39, 40 - 47 and 48 - 55.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 39.
-Low connectivity plot for the PolarStar graph for the length $l$ = 5.
+Figure 39: Low connectivity plot for the PolarStar graph for the length $l$ = 5.
 The structure graph is an ER2 graph and the subgraph is a BDF3 graph.
 We can clearly see, that for $l$ = 5 this PolarStar graph has a good connectivity $\geq$ 75% for all router pairs.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 40.
-Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 3.
+Figure 40: Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 3.
 Each cluster has a good connectivity within itself.
@@
 The connectivity between the two clusters is very low and only about half of the nodes in each cluster have a path to nodes in the other cluster.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 41.
-Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 5.
+Figure 41: Low connectivity plot for the Cascade Dragonfly topology with two clusters for the length $l$ = 5.
 Right underneath the first cluster, we can see the four node pairs that connect the two clusters.
 Even for paths of $l$ = 5, the connectivity between the two clusters stays low at only 10%.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 42.
-Low connectivity plot for the Tofu topology with X = Y = Z = 2 and $l$ = 3.
+Figure 42: Low connectivity plot for the Tofu topology with X = Y = Z = 2 and $l$ = 3.
 Not all nodes are connected with each other at length $l$ = 3.
@@
 This is the number of connections each node has to other clusters.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 43. Low connectivity plot for the Tofu topology with X = Y = Z = 2 and length $l$ = 6. There is a good connectivity $\geq$ 70% for most node pairs.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 44.
-Interference plots for Kautz and Arrangement graphs of length $l$ = 5.
+Figure 43: Low connectivity plot for the Tofu topology with X = Y = Z = 2 and length $l$ = 6. There is a good connectivity $\geq$ 70% for most node pairs.
+Figure 44: Interference plots for Kautz and Arrangement graphs of length $l$ = 5.
 The Arrangement networks have the same or less interference than their Jellyfish equivalent counterpart for $l\leq$ 5.
@@
 The same can be seen for $K_{4,n}$ and $K_{5,n}$ Kautz graphs, but not for $K_{6,n}$, which has the most interference for $l$ = 3 and $l$ = 4, but the least interference for $l$ = 5.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 45.
-Two-dimensional detailed interference plots for Kautz and Arrangement graphs of length $l$ up to 5.
+Figure 45: Two-dimensional detailed interference plots for Kautz and Arrangement graphs of length $l$ up to 5.
 $A_{n,4}$ has more interference than $A_{n,5}$.
@@
 Their equivalent Jellyfish have similar interference analysis, but in general a bit less.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 46.
-Interference plots for meshes of different dimensions of lengths up to $l$ = 5.
+Figure 46: Interference plots for meshes of different dimensions of lengths up to $l$ = 5.
 While 4D and 6D meshes have a pretty similar interference analysis, the 5D mesh has more interference.
@@
 For $l$ = 5, all meshes have more node pairs with zero interference compared to their equivalent Jellyfish graphs.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 47.
-Two-dimensional detailed interference plots for meshes of different dimensions of lengths up to $l$ = 5.
+Figure 47: Two-dimensional detailed interference plots for meshes of different dimensions of lengths up to $l$ = 5.
 We can see, that all meshes have similar interferences analysis.
 For the base mesh networks and $l$ = 5, all components are affected by interference, while for the equivalent Jellyfish, only a subset of them are.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 48.
-Interference plots for Tofu, 6D mesh and 6D torus with lengths up to $l$ = 5.
+Figure 48: Interference plots for Tofu, 6D mesh and 6D torus with lengths up to $l$ = 5.
 The Tofu topology has about the same amount of interference as the 6D torus network, but more than the regular 6D mesh.
 For $l$ = 5, both torus and the Tofu graph have more interference than the 6D mesh and more than any of the Jellyfish equivalents.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 49.
-Interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l$ = 1 up to $l$ = 5.
+Figure 49: Interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l$ = 1 up to $l$ = 5.
 Slim Fly has the most interference at $l$ = 3, but less interference for all other values of $l$.
@@
 The interference analysis looks similar for the topologies when comparing them to their equivalent Jellyfish, although all of the base topologies have slightly more interference.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 50.
-Two-dimensional detailed interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l\leq$ 5.
+Figure 50: Two-dimensional detailed interference plots for PolarStar topology with both subgraphs, Dragonfly and Slim Fly of length $l\leq$ 5.
 The detailed analysis of both PolarStar graphs are fairly similar.
 While for Dragonfly, the interference for $l$ = 4 is contributed by a lot of components, only a small amount of components contribute to the high amount of interference for PolarStar.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 51.
-Interference plots for Cascade Dragonfly, Spectralfly and PolarStar with both subgraphs of length $l$ = 5.
+Figure 51: Interference plots for Cascade Dragonfly, Spectralfly and PolarStar with both subgraphs of length $l$ = 5.
 Spectralfly has little interference compared to the other topologies for all lengths.
@@
 The most extreme difference when comparing to its equivalent Jellyfish, can be seen for $l$ = 4 and $l$ = 5.
-While the Jellyfish graph has at most interference 3 for any node quadruple, there are groups of nodes with interference values $>$ 15.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 52.
-Interference plots for Megafly with different gap sizes and Dragonfly of length $\leq$ 5.
+While the Jellyfish graph has at most interference 3 for any node quadruple, there are groups of nodes with interference values >> 15.
+Figure 52: Interference plots for Megafly with different gap sizes and Dragonfly of length $\leq$ 5.
 In general, Megafly has less interference than Dragonfly for length $l\leq$ 3.
@@
 This cannot be seen in the Dragonfly topology.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 53.
-Two-dimensional detailed interference plots for Megafly with different gap sizes and Dragonfly.
+Figure 53: Two-dimensional detailed interference plots for Megafly with different gap sizes and Dragonfly.
 For lengths of 2 and 3, only a few components contribute to the total interference in the Megafly networks.
@@
 This is also true for all Jellyfish equivalent graphs.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 54.
-Interference plots for k-ary n-trees for different values for k, Fat Trees and Fat Trees2x.
+Figure 54: Interference plots for k-ary n-trees for different values for k, Fat Trees and Fat Trees2x.
 For all base topologies, most node pairs have an interference of zero.
@@
 There are also more outlying node pairs than in the Fat Tree variants.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 55.
-Two-dimensional detailed interference plots for the k-ary 3-tree, Fat Tree and Fat Tree2x.
+Figure 55: Two-dimensional detailed interference plots for the k-ary 3-tree, Fat Tree and Fat Tree2x.
 For both components, there exist only specific values they can get.
 Since values in the diagonal cancel each other out, we can see that there is only one peak next to interference zero for all base topologies.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
-Figure 56.
-Interference plots for xGFTs of different variants and k-ary n-tree.
+Figure 56: Interference plots for xGFTs of different variants and k-ary n-tree.
 Similar to before, all topologies have interference of zero for most node pairs.
@@
 It has only interference higher than zero for $l>$ 3 and the peaks have only interference of 3.
-xGFT8 has only an interference $>$ 0 at length 2 and 3.
-[…]
-[Main findings are in the caption.]Main findings are in the caption.
+xGFT8 has only an interference >> 0 at length 2 and 3.